### v1

# Representational Analysis of Binding in Large Language Models

Abstract
Entity tracking is essential for complex reasoning. To perform in-context entity tracking, language models (LMs) must bind an entity to its attribute (e.g., bind a container to its content) to recall attribute for a given entity. For example, given a context mentioning “The coffee is in Box Z, the stone is in Box M, the map is in Box H”, to infer “Box Z contains the coffee” later, LMs must bind “Box Z” to “coffee”. To explain the binding behaviour of LMs, Feng and Steinhardt (2023) introduce a Binding ID mechanism and state that LMs use a abstract concept called Binding ID (BI) to internally mark entity-attribute pairs. However, they have not directly captured the BI determinant information from entity activations. In this work, we provide a novel view of the Binding ID mechanism by localizing the prototype of BI information. Specifically, we discover that there exists a low-rank subspace in the hidden state (or activation) of LMs, that primarily encodes the order of entity and attribute and which is used as the prototype of BI to causally determine the binding. To identify this subspace, we choose principle component analysis as our first attempt and it is empirically proven to be effective. Moreover, we also discover that when editing representations along directions in the subspace, LMs tend to bind a given entity to other attributes accordingly. For example, by patching activations along the BI encoding direction we can make the LM to infer “Box Z contains the stone” and “Box Z contains the map”. $\ast$ Equal contribution
[…]
Representational Analysis of Binding in Large Language Models
Qin Dai1, Benjamin Heinzerling2,1, Kentaro Inui3,1,2 1Tohoku University   2RIKEN AIP   3MBZUAI qin.dai.b8@tohoku.ac.jp, benjamin.heinzerling@riken.jp kentaro.inui@mbzuai.ac.ae
1 Introduction
The ability of a model to track and maintain information associated with an entity in a context is essential for complex reasoning Karttunen (1976); Heim (1983); Nieuwland and Van Berkum (2006); Barzilay and Lapata (2008); Kamp et al. (2010). To recall attribute information in the context, the model must bind entities to their attributesFeng and Steinhardt (2023). For example, given Sample 1 and 1, a model must bind the entities (e.g., “Box Z”, “Box M”, “Box H”, “Alex”, “John” and “Carl”) to their attributes (e.g., “coffee”, “stone”, “map”, “bean”, “pie” and “fruit”) respectively so as to accurately recall such as what is in “Box Z” or what is sold by “Alex” without confusion. Binding has also been studied as a fundamental problem in Psychology Treisman (1996). To uncover how Language Models (LMs) realize binding in term of internal representation, Feng and Steinhardt (2023) introduce a Binding ID mechanism and state that LMs apply a abstract concept called Binding ID (BI) to bind and mark Entity-Attribute (EA) pairs (e.g., “Box Z” and “coffee”, as shown in Sample 1, where BI is denoted as a numbered square). They also claim that the BI is represented as a vector to be added on the representation (or activation) of a EA pair so that the common vector is used as a key clue to search attribute for a given entity. However, they have not directly captured the prototype of BI information (called BI information hereafter for simplicity) from the activations.
[…]
Figure 1: Our main finding on Binding ID (BI) subspace intervention. Patching entity (e.g., "Z") representations along BI direction in activation space yields corresponding changes in model output.
Since binding is the foundational skill that underlies entity tracking Feng and Steinhardt (2023), in this work, we take the entity tracking task Kim and Schuster (2023); Prakash et al. (2024) as a benchmark to evaluate the LM’s binding behaviour. Based on the analysis of internal representation on this task, we localize the BI information from the activiations and provide a novel view of the Binding ID mechanism. Specifically, we discover that LMs encode (or store) BI information into a low-rank subspace (called BI subspace hereafter) that primarily encodes the order (i.e., the input order from left to right) of entity and attributes and which is used as the prototype of BI to determine binding. To identify the BI subspace, we take Principle Component Analysis (PCA) as our first attempt  [^1] Besids PCA, we also attempt partial least squares regression for capturing BI subspace. Since they achieve similar regression score, we adopt PCA for simplicity. See Appendix (§A.2) for details. to capture the subspace representing the BI information, and it is empirically proven be effective. Therefore, our findings confirm and extend prior understanding of BI in LMs.
Going beyond the prior work on BI mechanism Feng and Steinhardt (2023), we show that by causally intervening along the BI encoding Principle Component (PC), LMs swap the binding and infer a new attribute for a given entity accordingly. That is, we find a consistent causal relationship between the BI subspace intervention and the inferred attributes by LMs. For example, as shown in Figue 1, by patching activations along a direction (i.e., PC1), we can make the LMs to infer "Box Z contains the stone" and "Box Z contains the map" instead of "Box Z contains the coffee".
Overall, our findings suggest that LMs encodes BI information into a subspace of entity activations that primarily encodes the appearance order (or reversed order) of a EA pair in a given context. Moreover, the discovered BI subspace plays a crucial role in the in-context binding computation. In addition, we find that BI subspace not only exists in the Pretrained large LMs such as Llama2 Touvron et al. (2023) and Llama3 AI@Meta (2024), but also in the code fine-tuned LM such Float-7B Prakash et al. (2024).
2 Finding Low-rank Subspace for Binding ID
In this section we describe our Principle Component Analysis (PCA) based method to locate the BI subspace in activations of LMs. Firstly we extract entity activation from LMs as shown in Figure 1. Given a LM (e.g., Llama2), and a collection of texts which describe a set of EA pairs for a relation such as Sentence 1 for relation “is_in”, we extract the activation of entity token (e.g., “Z”) in query (denoted as $\mathbf{x}_{i}$) from certain layer  [^2] See Appendix (§A.3) for the layer selection. and construct a activation matrix $M_{r}\in R^{n\times d}$ for a relation $r$, where $n$ denotes the number of entities and $d$ denotes the dimension of the activation. The row $i$ of $M_{r}$ is the activation of an entity token (i.e., $\mathbf{x}_{i}$).
PCA has been applied for identifying various subspace (or direction) such as the subspace encoding language bias Yang et al. (2021), truth value of assertions Marks and Tegmark (2023) and sentiment Tigges et al. (2023). Inspired by these studies, we choose PCA as our first attempt to localize BI subspace. We hypothesize that in a activation subspace, entities with the same BI tend to cluster together (w.r.t the ones with different BIs), even though these entities have different semantic meaning, and the BIs are encoded as directions (or a PC) in the subspace. For convenience, we number BIs in left-to-right order, and the leftmost BI$=0$.
To capture the subspace, or BI direction, we leverage PCA, which identifies the principle directions of a space. Specifically, the PCA of a activation matrix is ${M_{r}}=U_{r}\Sigma_{r}V^{T}_{r}$, where the columns of $V_{r}\in R^{d\times d}$ are principle directions of $M_{r}$. We takes first $c$ columns of $V_{r}$ as the BI direction, denoted as $B_{r}\in R^{d\times c}$.
3 BI Subspace Visualization
Figure 2: Layer-wise BI subspace visualization on Llama2-7B.
We adopt a subset of the entity tracking dataset Kim and Schuster (2023); Prakash et al. (2024), which contains $n=1000$ samples, to create layer-wise activation matrix $M_{r}^{l}$. We then uses the $M_{r}^{l}$ to extract the layer-wise BI subspace projection matrix $B_{r}^{l}\in R^{d\times 2}$ to visualize the activation. Figure 2 shows the embedding visualization on Llama2-7B, where each point represents the activation of an entity projected via the $B_{r}^{l}$, and the colors represent BIs. From which, we can observe that middle layers, such as layer 8, have a clearly visible direction along which BI increases (or decreases), while the others have tangled distribution. We also observe similar property of distribution on Llama3-8B and Float-7B (§A.4). This indicates that LMs use the middle layers to encode BI information. We call this dimension that represents the order of BI as BI Principle Component (BI-PC). In the following section, we apply causal intervention on the BI-PC to analyze how BI-PC affect the model output.
4 Causal Interventions on Binding ID Subspace
Answer for # Step
Context Query 1 2 3 4 5 6
The coffee is in Box Z, the stone is in Box M, the map is in Box H, the coat is in Box L, the string is in Box T, the watch is in Box E, the meat is in Box F. Box Z contains the stone map map string watch meat
The letter is in Box Q, the boot is in Box C, the fan is in Box N, the crown is in Box R, the guitar is in Box E, the bag is in Box D, the watch is in Box K. Box Q contains the boot fan crown guitar watch watch
The cross is in Box Z, the ice is in Box D, the ring is in Box F, the plane is in Box Q, the clock is in Box X, the paper is in Box I, the engine is in Box K. Box Z contains the ice ring ring clock paper engine
Table 1: Attributes inferred by Llama2-7B as a result of directed activation patching along BI-PC in the BI subspace on the dataset of “r: is_in”, where color denotes the BI.
By projecting the activation matrix $M_{r}$ into the BI subspace, we have found a correlative evidence for the existence of the direction (i.e., BI-PC) that encodes Binding IDs. However, it is possible that the BI information is encoded in the BI subspace but has no effect on model output.
In order to test if Binding IDs are not only encoded in the BI subspace, but that these representations can be steered so as to swap the binding and change LM’s output. We now perform interventions to establish the causality. That is, we want to find out if making BI swapping interventions leads to a change in model output.
4.1 Activation Patching
Since LM computation graph could be viewed as causal graph Meng et al. (2022); McGrath et al. (2023), we intervene on model activations via activation patching Vig et al. (2020); Wang et al. (2022); Zhang and Nanda (2023); Heinzerling and Inui (2024); Engels et al. (2024), and observe the effect on model output. Unlike the common activation patching setup in which one replaces activations resulting from an original input with activations from a corrupted one, we create patches by editing activations along a particular direction (i.e., along BI-PC), similar to the activation editing method of Matsumoto et al. (2023); Heinzerling and Inui (2024); Engels et al. (2024). Although automatic methods for localizing model circuits of interest have been proposed Conmy et al. (2023); Kramár et al. (2024), for simplicity we perform a coarse layer-wise search based on the effect of activation patching in a development set, as shown in Appendix (§A.3), and use the found setting for all experiments.
4.2 Setting
Dataset To explore the internal representation that enables binding, we adopt the entity tracking dataset Kim and Schuster (2023); Prakash et al. (2024). The dataset contains English sentence describing a set of objects located in a set of boxes with difference labels, and the task is to infer what is inside a given box. For instance, when a LM is presented with “The coffee is in Box Z, the stone is in Box M, the map is in Box H, … Box Z contains the”, it should infer the next token as “coffee”. Each sample involves $7$ AE pairs. To evaluate the binding in various context, we also apply the templates shown in Table 2 to generate other $5$ datasets with different relation, where $a_{i}$ and $e_{i}$ denotes the attribute and entity, and they are sampled from a fixed pool of $224$ one-token objects (e.g., “dog”, “corn” and “cookie”) and $523$ of one-token names (e.g., “Alex”, “Juli” and “Dan”) respectively. We sample $n=1000$ context from each dataset to run the following analysis.
Template
$1$ The $a_{0}$ is sold by person $e_{0}$, …, the $a_{i}$ is …, $a_{7}$ is sold by person $e_{7}$. Person $e_{i}$ is selling the
$2$ The $a_{0}$ is applied by person $e_{0}$, …, the $a_{i}$ is …, $a_{7}$ is applied by person $e_{7}$. Person $e_{i}$ applies the
$3$ The $a_{0}$ is moved by person $e_{0}$, …, the $a_{i}$ is …, $a_{7}$ is moved by person $e_{7}$. Person $e_{i}$ moved the
$4$ The $a_{0}$ is brought by person $e_{0}$, …, the $a_{i}$ is …, $a_{7}$ is moved by person $e_{7}$. Person $e_{i}$ brings the
$5$ The $a_{0}$ is pushed by person $e_{0}$, …, the $a_{i}$ is …, $a_{7}$ is pushed by person $e_{7}$. Person $e_{i}$ pushes the
Table 2: Templates of Dataset.
Metrics We apply two evaluation metrics. The logit difference metric introduced in Wang et al. (2022), which calculates difference in logits between original and intervened answers, as well as the "logit flip" accuracy metric Geiger et al. (2022), which represents the proportion of cases where we alter the model output after a causal intervention.
(a) r: is_in
(b) r: sell
(c) r: apply
(d) r: move
(e) r:bring
(f) r: push
Figure 3: Logit Difference (LD) for BI-PC based intervention across datasets on Llama2-7B, where x axis denotes the number of intervention steps on $e_{0}$, y axis does the LD, BI_i represents each target attribute and the light yellow bottom line indicates the LD of original attribute (i.e., $a_{0}$). Here, $l=8$, $v=2.5$, and $\alpha=3.0$.
4.3 Results: Direct Editing BI Subspace
We hypothesize that LMs encode BI information into a low-rank BI subspace. Therefore, we wonder if a LM changes the binding behavior, when adding a particular value $v$ (called step hereafter) along the BI-PC mentioned in Section (§2). For example, if we add one unit of $v$ on the activation of $e_{0}$, the LM will reset its BI as $1$ and bind attribute $a_{1}$ to entity so that infers the $a_{1}$ as the attribute of $e_{0}$ instead of the original $a_{0}$. Similarly, adding two units of $v$ will make the LM infer $a_{2}$ for $e_{0}$, and so on. We intervene via the Equation 1, where $\mathbf{x}_{0,l}$ is the original activation of $e_{0}$ (i.e., the leftmost entity) in layer $l$, $\mathbf{x}_{0,l}^{*}$ is the intervened activation, $B_{r}$ is the BI subspace projection matrix mentioned in Section (§2), $\alpha$ is a hyper-parameter to scale the effect of intervention and $\beta$ ($0\leq\beta\leq 6$) denotes the number of steps.
$\mathbf{x}_{0,l}^{*}=\mathbf{x}_{0,l}+\alpha B_{r}^{T}(B_{r}\mathbf{x}_{0,l}+\beta v)$ (1)
Table 1 lists several examples under the BI subspace intervention on the entity tracking dataset Kim and Schuster (2023); Prakash et al. (2024). We also list the examples from other datasets in Appendix (§A.6). We can see that when adding $1$ step along BI-PC, the model selects “stone” for entity “Z” instead of its original attribute “coffee”. Similarly, when the step is doubled, the model will select attribute “map” for the entity, and so on. Although the attribute selection does not strictly follow the number of steps, this indicates, to some extent, that changing the value along BI-PC can induce the swap of attribute.
Besides the qualitative analysis, we also conduct quantitative analysis for the causality between the BI subspace based activation patching and the binding behaviour of the LM. We plot mean-aggregated effect of directed activation patching across multiple datasets in Figure 3. Figure 3 indicates how the Logit Difference (LD) of each attribute changes as the step increases. We can observe that as the number of steps increases, LD of the original attribute decreases. In contrast, LD of other attributes gradually increase until a certain point and then gradually decrease. Given a candidate attribute, its peak point generally corresponds the number of steps that equals to its BI. For instance, when adding $3$ steps, the points of BI_3 (i.e., attributes of BI$=3$) on step$=3$ achieve the highest LD score. This indicates that by adjusting the value along the BI-PC, we can increase the probability of the corresponding attribute, thereby swap the answer.
Similarly, Figure 4 illustrates the relation between the number of steps and the logit flip, which gauges the percentage of the predicted attributes under an intervention. Figure 4 shows that as the step increases, the proportion bar becomes darker, it means that the model promotes the proportion of the corresponding following attribute in its inference. For instance, when adding $3$ step on the subspace, the $a_{3}$ (i.e., BI_3) becomes the major of the answers. This proves that the discovered subspace stores BI information, and the subspace will causally affect the computation of Binding in a LM. See Appendix (§A.9) for the results on Llama3-8B.
(a) r: is_in
(b) r: sell
(c) r: apply
(d) r: move
(e) r: bring
(f) r: push
Figure 4: Logit flip for BI-PC based intervention across datasets on Llama2-7B, where x axis denotes the number of intervention steps on $e_{0}$, y axis does the proportion of each inferred attribute in model output.
4.4 Results: Activation Steering on BI Subspace
Figure 5: Logit Difference and Logit Flip for activation steering on the entity tracking dataset (i.e., r: is_in), where x axis represents the intervention of $\mathbf{s}_{0\rightarrow bi}$.
Inspired by the research on Activation Steering (AS) Turner et al. (2023), we apply an AS method to verify the importance of the BI subspace on LM’s binding behaviour. Specifically, we use the following Equation 2 to extract a subspace steering vector $\mathbf{s}_{0\rightarrow bi}$, which is proposed to swap BI from 00 to $bi$, where $n$ is the number of target entities, $\mathbf{x}_{bi,l}^{i}$ represents the activation of entity $e_{i}$ from layer $l$, and its BI is $bi$. We intervene via Equation 3 and assume that by adding $\mathbf{s}_{0\rightarrow bi}$ to the original activation $\mathbf{x}_{0,l}$, we can increase the LD and the proportion of the attribute $a_{bi}$. Figure 5 shows the results on the entity tracking dataset Kim and Schuster (2023); Prakash et al. (2024). (Appendix (§A.7) shows the results on other datasets) These results indicate that AS can achieve the similar tendency as the direct value intervention mentioned in Section (§4.3). For instance, adding $\mathbf{s}_{0\rightarrow 3}$, which is used to swap BI form 00 to $3$, can increase the LD of $a_{3}$ and its proportion in the predicted answers. The consistent tendency with the results of the direct subspace editing, shown Figure 3 and Figure 4, further illustrates that the discovered subspace contains the BI information, and more importantly, it plays an important role when the model performs in-context binding computation.
$\mathbf{s}_{0\rightarrow bi}=\frac{1}{n}\sum_{i=1}^{n}(B_{r}\mathbf{x}_{bi,l}^{i}-B_{r}\mathbf{x}_{0,l}^{i})$ (2)
$\mathbf{x}_{0,l}^{*}=\mathbf{x}_{0,l}+\alpha B_{r}^{T}\mathbf{s}_{0\rightarrow bi}$ (3)
4.5 Binding Subspace and Position
Input (original)
$C_{1}$ $a_{0}^{p0}\,r\,e_{0}^{p1},\,a_{1}^{p2}\,r\,e_{1}^{p3},\,a_{2}^{p4}\,r\,e_{2}^{p5}.\quad e_{1}^{p3}\ r^{-1}\ ?$
$C_{2}$ $a_{0}^{p0}\,r\,e_{0}^{p1},\,a_{1}^{p2}\,r\,e_{1}^{p3},\,a_{2}^{p4}\,r\,e_{2}^{p5}.\quad e_{2}^{p5}\ r^{-1}\ ?$
Input (with pseudo)
$C^{\prime}_{1}$ $a_{*0}^{p0}\,r\,e_{*0}^{p1},\,a_{*0}^{p2}\,r\,e_{*0}^{p3},\,a_{1}^{p4}\,r\,e_{1}^{p5}.\quad e_{1}^{p5}\ r^{-1}\ ?$
$C^{\prime}_{2}$ $a_{*0}^{p0}\,r\,e_{*0}^{p1},\,a_{1}^{p2}\,r\,e_{1}^{p3},\,a_{2}^{p4}\,r\,e_{2}^{p5}.\quad e_{2}^{p5}\ r^{-1}\ ?$
Table 3: Simplified expression of original inputs and the one modified with pseudo relation, which is proposed to equalize PI for PCA analysis, where $a_{0}^{p1}\,r\,e_{0}^{p2}$ represents a relation such as “the apple is in Box C”, and $e_{0}^{p2}$ denotes an entity with BI of 00 and PI of $p2$, $e_{2}^{p5}\ r^{-1}\ ?$ denotes the query on entity $e_{1}$, such as “Box C contains the”.
Input (original)
$a_{0}^{p0}\,r\,e_{0}^{p1},\,a_{1}^{p2}\,r\,e_{1}^{p3},\,a_{2}^{p4}\,r\,e_{2}^{p5}.\quad e_{0}^{p1}\ r^{-1}\ ?$
Input (with interjection)
$a_{0}^{p0}\,r\,e_{0}^{p1},\,j^{p2}\,j\,j^{p3}\,...j^{pi}\,a_{1}^{pi+1}\,r\,e_{1}^{pi+1}...\quad e_{0}^{p1}\ r^{-1}\ ?$
Table 4: Simplified expression of original input and the one modified with a sequence of interjections, where $j^{p3}$ denotes an interjection $j$, such as “ah”, with position of $p3$, which is also the position of $e_{1}^{p3}$ in the original input.
Input (original)
“The apple is in Box E, the bell is in Box F, …”
Input (with filler words)
I will find out that the apple is in Box E, the bell is in Box F, …“
Table 5: An example of the dataset with filler words “I will find out that”.
In this section, we discuss the relationship between the BI subspace and Positional Information (PI), namely the $postion\_ids$ of input tokens. As mentioned in Section (§4.3), the discovered subspace stores the BI information, therefore direct intervention on it can swap the answer of LM. However, one counter hypothesis is that the subspace is not used for storing BI information but the PI of attributes, thus the direct intervention merely change the PI so that swap the answer. Regarding the relationship between BI and PI, Feng and Steinhardt (2023) found that even when PI of attributes is swapped, the model still makes correct predictions, thus confirming the independence between BI and PI. Based on this finding, we go one step further and illustrate the independence between the BI subspace and PI. To prove the independence, we create three datasets, the first one is by extending the original dataset with pseudo relation, as shown in Table 3, the second one is by inserting a sequence of interjection(e.g., “ah, ah, …”), as listed in Table 4, and the third one is by prefixing the original dataset with filler words(e.g., “It can be seen that”).
New Dataset with Pseudo Relation. In Table 3, $a_{*0}^{p0}\,r\,e_{*0}^{p1}$ refers to a pseudo relation, which is a fixed expression, such as “the PC is in Box Z”, applied to adjust the PI while keeping the BI. For instance, in Table 3, adding one or two $a_{*0}^{p0}\,r\,e_{*0}^{p1}$ before $a_{1}\,r\,e_{1}$ (i.e., $C^{\prime}_{2}$ and $C^{\prime}_{1}$) does not affect the BI of $e_{1}$ but its PI, because $e_{1}$ is still the second unique entity from the left (i.e., BI$=1$), but its PI is $p3$ and $p5$ respectively. Using the pseudo relation, we create the data in a manner that the target entity (e.g., $e_{1}^{p5}$ and $e_{2}^{p5}$) to extract activation have the same PI, such as $C^{\prime}_{1}$ and $C^{\prime}_{2}$ in Table 3.
We apply the method mentioned in Section (§2) on the set of activations $\{\overrightarrow{e_{1}^{p5}},\overrightarrow{e_{2}^{p5}},...\}$, where $\overrightarrow{e_{1}^{p5}}$ denotes the activation of $e_{1}$ in $e_{1}^{p5}\ r^{-1}\ ?$, so as to capture the BI difference and exclude the PI difference, because they share the same PI (i.e., $P5$) but different BI (i.e., $1$, $2$, …). Then we compare its BI subspace with the original one (e.g., $\{\overrightarrow{e_{1}^{p2}},\overrightarrow{e_{2}^{p5}},...\}$) to analyze how the distribution of BI subspace changes after removing the PI variance. Figure 6 visualizes the BI subspace distribution, where the light colored points denote the original distribution, and the dark ones are from the new one with equalized PI. We can observe that after removing the PI difference, the distribution is still similar to the original one that there is a clearly visible direction along which BI increases. This illustrates that our PCA based method can capture BI information, that is, along the direction of BI-PC, and it does not causally depend on PI. See Appendix (§A.5) for our further analysis on how the context of binding affects the distribution.
New Dataset with Interjections. The dataset is created by inserting a sequence of interjections after the first attribute entity pair, as illustrated in Table 4. Since there is no BI information in the interjection (e.g., $j^{p3}$), adding it only changes the PI of its following entities and attributes. We set the number of interjections as that its last PI is larger than the last PI of its original input (e.g., $pi>p5$ in Table 4).
Based on this dataset, we conduct the same intervention on its BI subspace, as mentioned in Section (§4.3). The counter argument is that the subspace only captures PI, and the intervening step only changes the PI information. Specifically, adding one unit of $v$ on $\overrightarrow{e_{0}^{p1}}$ can change its PI from $p1$ to $p3$, which is the PI of $e_{1}$, and its attribute is $a_{1}$, as shown in Table 4, thereby the LM swaps the answer from $a_{0}$ to $a_{1}$. If it is true, then the same intervention will not change the answer on the new dataset, because following the counter argument, after adding one unit of $v$ on $\overrightarrow{e_{0}^{p1}}$, its PI becomes $p3$, and $p3$ is the PI of $j^{p3}$. The LM thus would not select $a_{1}$ as its answer. However, the results on Figure 7 shows that the subspace intervention on the new dataset achieves similar results as the original one, as shown in Figure 3 and Figure 4, proving the counter argument wrong and indicating the independence between BI subspace and PI.
New Dataset with Filler Words. The dataset is created by adding Filler Words(FW) with various length, such as “OK”, “I see that” and “There is no particular reason”, in front of the entity tracking dataset Kim and Schuster (2023); Prakash et al. (2024), as shown in Table 5. Since the length (i.e., the number of tokens) of FW directly affects the PI of its following entities and attributes without affecting their BI, we take the length as the measure of intervention on PI and apply Spearman’s rank correlation $\rho$ to calculate the correlation between the length (denoted as PI) and the discovered BI-PC. Figure 8 shows $\rho$ between PI and BI-PC as well as between BI and BI-PC. We can observe that BI-PC has high $\rho$ with BI but almost zero $\rho$ with PI, indicating that the discovered BI-PC is dependent on BI but independent on PI. Therefore, the BI-PC does not simply encode absolute token position.
Figure 6: Embedding visualization for activation with equalized PI, where “Ori” denotes the distribution of original dataset, while “Pse” denotes the distribution of the new dataset with pseudo relation.
Figure 7: Logit Difference and Logit Flip for activation patching on the dataset of “r: is_in” with interjections. Appendix (§A.10) shows the results on other dataset with the same interjection based modification.
[…]
4.6 Binding Subspace and Relational Information
Input (original)
“The apple is in Box E, the bell is in Box F, …”
Input (Non-related)
“I see apple, cup and hair, somewhere else there is Box E, the bell, bus, toy, cookie and Box F are scattered around, ”
Table 6: An example of the dataset with non-related expression.
In order to study the relationship between BI-PC and Relational Information(RI), we create an alternative dataset by converting relational expression of the entity tracking dataset Kim and Schuster (2023); Prakash et al. (2024) into non-related one. Specifically, we prepare a set of non-related expression templates and randomly select one to replace the original expression of relation as shown in Table 6. We can observe that the template could make a target EA pair (e.g., “apple” and “Box E”) semantically unrelated, but retain their Ordering Information(OI) (i.e., “apple”->“bell”, and “Box E”->“Box F”).
The BI mechanism Feng and Steinhardt (2023) notes that a LLM represents binding through the consistency of BI. According to the mechanism, we adopt the correlation score between BI-PC of attributes (e.g., BI-PC of “apple”) and of their corresponding entities (e.g., BI-PC of“Box E”) to represent their binding strength. We select Spearman’s rank correlation $\rho$ as the correlation score and present the $\rho$ of the non-related dataset as well as the related one in the Figure 9.
We can observe that there is strong correlation between BI-PC of attributes and of their corresponding entities in the original data, indicating that LLM applies the correlation of BI-PC to encode binding. We can also observe that $\rho$ of non-related dataset is lower than the related (i.e., original) one, indicating that the BI-PC contains a certain degree of RI so that removing RI can decrease the $\rho$. However, even there is no RI but OI in the non-related dataset, the BI-PC of attributes still has high correlation with their co-occurring entities, indicating that BI-PC not only stores RI, but also encodes OI, and the latter seems to be the main and fundamental factor that is utilized by the LLM to determine binding, because intervention on BI-PC can causally change the behaviour of LLM (as discussed in Section (§4). Therefore, we hypothesize that the BI-PC primarily encodes the OI of entities and attributes that is used as a prototype of BI to causally determine binding.
[…]
4.7 Results on Fine-Tuned LM
Kramár et al. (2024); Kim et al. (2024) claim that the code fine-tuned LM, such as Float-7B Prakash et al. (2024) outperforms the pretrained LM on the entity tracking task Kim and Schuster (2023); Prakash et al. (2024). Since the code fine-tuned LM performs well on the entity tracking task that requires the BI subspace based computation, we hypothesize that BI subspace also exists in the code fine-tuned LM and the intervention along BI-PC will causally affect the model output. To prove the hypothesis, we conduct the intervention on Float-7B and show results in Figure 10. We found that the BI subspace based intervention on Float-7B achieves the similar results as on Llama2-7B, indicating that the BI subspace not only exists in the pretrained LM but also in the fine-tuned one. In addition, adding the same step value (i.e., $v$) on Float-7B will achieve higher LD value than Llama2-7B, indicating that the code fine-tuned LM is more sensitive to the BI subspace based intervention. For instance, the maximum LD of $a_{4}$ in the former is around $10$, and it is $2$ times larger than the one in the latter, which is around $5$. This might partially explains why the code fine-tuned LM performs better than the original one, because code fine-tuning might enhance the function of BI subspace so that it is more sensitive on the intervention and more effective on the in-context entity tracking task.
Figure 10: Logit Difference and Logit Flip for activation patching on the entity tracking dataset (i.e., r: is_in) on Float-7B. See Appendix (§A.8) for the results on other datasets.
5 Related Work
Linear Representation Recent research found that sequence models trained only on next token prediction linearly represent various semantic concepts including Othello board positions Li et al. (2022); Nanda et al. (2023), the truth value of assertions Marks and Tegmark (2023), sentiment Tigges et al. (2023), and numeric values such as elevation, population, birth year, and death year Gurnee and Tegmark (2023); Heinzerling and Inui (2024). Continuing this line of research, in this work, we discover that LMs such as Llama-2 can also linearly encode BI, because there is a linear direction that primarily encodes BI in the activations. See Appendix (§A.1) for other related work.
6 Conclusion and Future Work
In this work, we study the in-context binding, a fundamental skill underlying many complex reasoning and natural language understanding tasks. We provide a novel view of the Binding ID mechanism introduced by Feng and Steinhardt (2023) that there exists a low-rank subspace in the hidden state (or activation) of LMs that primarily encodes the ordering information and which is used as the prototype of BIs to causally determine binding. Our future work includes: 1. the analysis of BI subspace in a more realistic setting; 2. the study of interaction between in-context binding and factual knowledge learned from pretraining; 3. BI subspace based mechanistic analysis.
Limitation
The limitations of our research include the following points. Firstly, we only analyze BI subspace on the attribute prediction task, but not on the entity inference task (i.e., given an attribute to infer its entity). Secondly, we lack the analysis on how predicate (or relation) affect the BI subspace, and how the results of BI-subspace based intervention differ with the type of predicate. Thirdly, although we use a publicly available entity tracking dataset, it is still a synthesized dataset. Therefore, for uncovering how LMs bind and track entity in reality, it is necessary to analyze the BI subspace on a real world dataset. The last but not lest, we only analyze binding from the perspective of representation and localize BI subspace. However, we have not answered what is the mechanism that generates the subspace and what is the circuit that utilizes the subspace for binding.
Ethical Statement
The existing dataset Kim and Schuster (2023); Prakash et al. (2024) and LMs (i.e., Llama2-7B, Float-7B and Llama3-8B) are applied according to their intended research purpose. The synthetic datasets we adopted in this work are automatically created by strictly following the rule (or pattern) of the existing dataset, where the entities and attributes are sampled from a pool of wide variety of one-token names and concepts. Therefore, there is no ethical concern on human annotation bias and semantic biases. The datasets and code will be publicly available to ensure the reproducibility of our experiments.
Appendix A Appendix
A.1 Other Related Work
Knowledge Localization Many works aim to localize and edit factual relations (e.g., “capital of”) that LMs learn from pretraining and are stored into model weights Geva et al. (2020); Dai et al. (2021); Meng et al. (2022); Geva et al. (2023); Hernandez et al. (2023). Different from this line of research, this work studies in-context representations of relations and analyzes how they are stored in model activations.
Mechanistic Interpretability Notable progress has been made in uncovering circuits performing various tasks within LMs Elhage et al. (2021); Wang et al. (2022); Wu et al. (2024). Recently, Prakash et al. (2024) identify the circuit for entity tracking task. Feng and Steinhardt (2023) introduce a Binding ID Mechanism for explaining the binding problem, state that LMs use the abstract concept BI to internally mark entity-attribute pairs. However, they does not directly capture BI information from activations. Therefore, they have not answered how LMs store the BI information into entity activations, how to localize the BI information and whether the localized BI information causally affect the model binding behaviour.
A.2 Partial least squares regression and PCA
Besides PCA, a commonly used unsupervised Dimension Reduction (DR) method, we also attempt Partial Least Squares regression (PLS) Wold et al. (2001), a supervised DR method. PLS extracts a set of ordered latent variables that maximizes the co-variability between the features (e.g., activations) and the scores to be predicted (e.g., BI). We perform PCA and PLS on a development set and compare their regression cures in Figure 11. We can observe that both the first PCA component and the first PLS direction contain almost all information about BI of target entity, because their regression score is close to one. The consistency indicates that PCA is an effective method to capture BI.
Figure 11: Regression curves for PLS and PCA.
A.3 Layer-wise Intervention
To determine how well BI subspace from different layers of a LM causally affects the model output, we perform layer-wise BI-PC based intervention mentioned in Section (§4.3) on our development set. In Figure 12, we can observe that BI subspace from middle layers (i.e., from layer7 to layer15, especially layer8) significantly affect the computation of binding, and interestingly, these layers also overlap with the ones that clearly encode BI information, as shown in Figure 2. Based on such analysis, we select the layer to perform activation patching.
Figure 12: Average Logit Difference (LD) and logit flip for layer-wise BI-PC based intervention on Llama2-7B, where x axis denotes the layer, the colored zone indicates the layers that are sensitive to the intervention, and the vertical line represents the most sensitive one (i.e., Layer 8), Y axis denotes the average LD and the proportion of inferred attributes (excluding the original one) in Figure 12(a) and Figure 12(b) respectively.
A.4 Layer-wise Embedding Visualization on Llama3-8B and Float-7B
(a) Llama3-8B
(b) Float-7B
Figure 13: Layer-wise BI subspace visualization on Llama3-8B and Float-7B.
A.5 Layer-wise Embedding Visualization after Masking Binding Information
The counter argument is that the captured subspace only represents the positional information. To test the claim, we attempt to mask the context around the entities and attributes with random two-letter tokens (e.g., “td”) so as to ablate the binding information and keep the positional information. For instance, we convert “the document is in Box Q , the bus is in Box F, …” as “pl document td cy wa Q br fl bus ti eq fs F …” so that there is no relational information in the latter. Figure 14 compares the BI subspace distribution between with and without binding information. We can observe that ablation of binding information tangles the distribution so that there is no clear clustering for each BI (e.g., by comparing layer14). This indicates that our discovered subspace encodes binding information.
(a) w/ Binding Information
(b) w/o Binding Information
Figure 14: Layer-wise BI subspace visualization for w/ and w/o binding information on Llama2-7B.
A.6 Case Study on Llama2-7B
Answer for # Step
Context Query 1 2 3 4 5 6
The bugis sold by person Esta, the spawn is sold by person Fritz, the wine is sold by person Inga, the paste is sold by person Ward, the poison is sold by person Albert, the crow is sold by person Davis, the nest is sold by person Val . Person Esta is selling the spawn wine paste poison crow nest
The virus is sold by person Anna, the fur is sold by person Earl, the pill is sold by person Flor, the bean is sold by person Roy, the spawn is sold by person Kam, the farm is sold by person Young, the sheep is sold by person Billy. Person Anna is selling the fur pill spawn spawn farm sheep
The root is sold by person Carl, the mouse is sold by person Marco, the fruit is sold by person Luke, the bug is sold by person Paul, the grass is sold by person Inga, the pie is sold by person Pok, the cookie is sold by person George. Person Carl is selling the mouse fruit bug grass cookie cookie
Table 7: Attributes inferred by Llama2-7B as a result of directed activation patching along BI-PC in the BI subspace on the dataset of “r: sell”, where color denotes the BI.
Answer for # Step
Context Query 1 2 3 4 5 6
The carbon is applied by person Wei, the liquid is applied by person Season, the bath is applied by person Robert, the fog is applied by person Daniel, the heavy is applied by person Roma, the motor is applied by person Ara, the pool is applied by person Jorge Person Wei applies the liquid bath fog motor motor pool
The rain is applied by person Kurt, the gauge is applied by person Jon, the dust is applied by person Newton, the jet is applied by person Dan, the floor is applied by person Alfred, the low is applied by person Mike, the basket is applied by person April Person Kurt applies the gauge dust jet floor basket basket
The lamp is applied by person Angel, the bucket is applied by person Carl, the canvas is applied by person Bert, the cargo is applied by person Otto, the plain is applied by person Johnny, the floor is applied by person John, the heavy is applied by person Era. Person Angel applies the bucket canvas cargo plain floor heavy
Table 8: Attributes inferred by Llama2-7B as a result of directed activation patching along BI-PC in the BI subspace on the dataset of “r: apply”, where color denotes the BI.
Answer for # Step
Context Query 1 2 3 4 5 6
The lip is moved by person Mack, the tract is moved by person Sommer, the pen is moved by person Son, the tip is moved by person August, the bat is moved by person Monte, the socket is moved by person Marco, the hook is moved by person Paul. Person Mack moved the tract pen tip bat hook hook
The mask is moved by person Jules, the timer is moved by person Ward, the bullet is moved by person Ana, the eye is moved by person Val, the button is moved by person Andy, the lock is moved by person Arnold, the colon is moved by person Betty. Person Jules moved the timer bullet button lock lock colon
The mask is moved by person Cole, the neck is moved by person Donald, the pad is moved by person Beth, the cone is moved by person Jorge, the tail is moved by person Lou, the thread is moved by person Alfred, the toe is moved by person Edward. Person Cole moved the neck pad cone tail toe toe
Table 9: Attributes inferred by Llama2-7B as a result of directed activation patching along BI-PC in the BI subspace on the dataset of “r: move”, where color denotes the BI.
Answer for # Step
Context Query 1 2 3 4 5 6
The creature is brought by person Tam, the guitar is brought by person Frank, the dress is brought by person Stuart, the block is brought by person Victor, the brain is brought by person David, the coffee is brought by person Mack, the radio is brought by person Roger. Person Tam brings the guitar dress block brain coffee radio
The boat is brought by person Luke, the pipe is brought by person Clara, the pot is brought by person Han, the bill is brought by person Chi, the milk is brought by person Scott, the card is brought by person Henry, the brick is brought by person Morris Person Luke brings the pipe pot bill card brick brick
The fan is brought by person Van, the note is brought by person Clara, the block is brought by person Alex, the newspaper is brought by person Peg, the crown is brought by person Jan, the car is brought by person Pok, the magnet is brought by person Golden. Person Van brings the note block crown car magnet magnet
Table 10: Attributes inferred by Llama2-7B as a result of directed activation patching along BI-PC in the BI subspace on the dataset of “r: bring”, where color denotes the BI.
Answer for # Step
Context Query 1 2 3 4 5 6
The load is pushed by person Mike, the atom is pushed by person Mira, the tin is pushed by person Juli, the stud is pushed by person Sam, the sedan is pushed by person Pia, the bath is pushed by person Leo, the growth is pushed by person Pat. Person Mike pushes the atom tin stud bath growth growth
The mud is pushed by person Thomas, the heavy is pushed by person Ralph, the tile is pushed by person Pierre, the import is pushed by person Perry, the arm is pushed by person Robert, the lung is pushed by person Kurt, the cabin is pushed by person Ernest. Person Thomas pushes the heavy tile import arm cabin cabin
The bed is pushed by person Fran, the lever is pushed by person Lan, the cord is pushed by person Paris, the vent is pushed by person Gene, the thumb is pushed by person Marie, the mouth is pushed by person Asia, the ear is pushed by person Lang. Person Fran pushes the lever cord vent thumb thumb ear
Table 11: Attributes inferred by Llama2-7B as a result of directed activation patching along BI-PC in the BI subspace on the dataset of “r: push”, where color denotes the BI.
A.7 Activation Steering on Llama2-7B
(a) r: is_in
(b) r: sell
(c) r: apply
(d) r: move
(e) r:bring
(f) r: push
Figure 15: Logit Difference (LD) for BI subspace based activation steering across datasets on Llama2-7B, where x axis represents the intervention of $\mathbf{s}_{0\rightarrow bi}$ on the activation of $e_{0}$. Here, $l=8$ and $\alpha=1.25$.
(a) r: is_in
(b) r: sell
(c) r: apply
(d) r: move
(e) r: bring
(f) r: push
Figure 16: Logit flip for BI subspace based activation steering across datasets on Llama2-7B, where x axis represents the intervention of $\mathbf{s}_{0\rightarrow bi}$ on the activation of $e_{0}$.
A.8 Activation Patching on Float-7B
(a) r: is_in
(b) r: sell
(c) r: apply
(d) r: move
(e) r:bring
(f) r: push
Figure 17: Logit Difference (LD) for BI-PC based intervention across datasets on Float-7B, where x axis denotes the number of intervention steps on $e_{0}$, y axis does the LD, BI_i represents each target attribute and the light yellow bottom line indicates the LD of original attribute (i.e., $a_{0}$). Here, $l=10$, $v=2.55$, and $\alpha=5.0$.
(a) r: is_in
(b) r: sell
(c) r: apply
(d) r: move
(e) r: bring
(f) r: push
Figure 18: Logit flip for BI-PC based intervention across datasets on Float-7B, where x axis denotes the number of intervention steps on $e_{0}$, y axis does the proportion of each inferred attribute in model output.
A.9 Activation Patching on Llama3-8B
(a) r: is_in
(b) r: sell
(c) r: apply
(d) r: move
(e) r:bring
(f) r: push
Figure 19: Logit Difference (LD) for BI-PC based intervention across datasets on Llama3-8B, where x axis denotes the number of intervention steps on $e_{0}$, y axis does the LD, BI_i represents each target attribute and the blue line indicates the LD of original attribute (i.e., $a_{0}$). Here, $l=10$, $v=0.65$, and $\alpha=2.0$.
(a) r: is_in
(b) r: sell
(c) r: apply
(d) r: move
(e) r: bring
(f) r: push
Figure 20: Logit flip for BI-PC based intervention across datasets on Llama3-8B, where x axis denotes the number of intervention steps on $e_{0}$, y axis does the proportion of each inferred attribute in model output.
A.10 Activation Patching on the New Dataset with Interjections
(a) r: is_in
(b) r: sell
(c) r: apply
(d) r: move
(e) r:bring
(f) r: push
Figure 21: Logit Difference for activation patching on the dataset with interjections. Here, $l=8$, $v=2.5$, and $\alpha=3.0$.
(a) r: is_in
(b) r: sell
(c) r: apply
(d) r: move
(e) r: bring
(f) r: push
Figure 22: Logit Flip for activation patching on the dataset with interjections.

### v3

@@
-Representational Analysis of Binding in Large Language Models
+Representational Analysis of Binding in Language Models
 Abstract
@@
 To explain the binding behaviour of LMs, Feng and Steinhardt (2023) introduce a Binding ID mechanism and state that LMs use a abstract concept called Binding ID (BI) to internally mark entity-attribute pairs.
-However, they have not directly captured the BI determinant information from entity activations.
-In this work, we provide a novel view of the Binding ID mechanism by localizing the prototype of BI information.
-Specifically, we discover that there exists a low-rank subspace in the hidden state (or activation) of LMs, that primarily encodes the order of entity and attribute and which is used as the prototype of BI to causally determine the binding.
-To identify this subspace, we choose principle component analysis as our first attempt and it is empirically proven to be effective.
-Moreover, we also discover that when editing representations along directions in the subspace, LMs tend to bind a given entity to other attributes accordingly.
-For example, by patching activations along the BI encoding direction we can make the LM to infer “Box Z contains the stone” and “Box Z contains the map”.
-$\ast$ Equal contribution
+However, they have not captured the Ordering ID (OI) from entity activations that directly determines the binding behaviour.
+In this work, we provide a novel view of the BI mechanism by localizing OI and proving the causality between OI and binding behaviour.
+Specifically, by leveraging dimension reduction methods (e.g., PCA), we discover that there exists a low-rank subspace in the activations of LMs, that primarily encodes the order (i.e., OI) of entity and attribute.
+Moreover, we also discover the causal effect of OI on binding that when editing representations along the OI encoding direction, LMs tend to bind a given entity to other attributes accordingly.
+For example, by patching activations along the OI encoding direction we can make the LM to infer “Box Z contains the stone” and “Box Z contains the map”.
+The code and datasets used in this paper are available at https://github.com/cl-tohoku/OI-Subspace.
 […]
-Representational Analysis of Binding in Large Language Models
+Representational Analysis of Binding in Language Models
 Qin Dai1, Benjamin Heinzerling2,1, Kentaro Inui3,1,2 1Tohoku University   2RIKEN AIP   3MBZUAI qin.dai.b8@tohoku.ac.jp, benjamin.heinzerling@riken.jp kentaro.inui@mbzuai.ac.ae
@@
 (2010).
-To recall attribute information in the context, the model must bind entities to their attributesFeng and Steinhardt (2023).
-For example, given Sample 1 and 1, a model must bind the entities (e.g., “Box Z”, “Box M”, “Box H”, “Alex”, “John” and “Carl”) to their attributes (e.g., “coffee”, “stone”, “map”, “bean”, “pie” and “fruit”) respectively so as to accurately recall such as what is in “Box Z” or what is sold by “Alex” without confusion.
+To recall attribute information for a given entity in a context, the model must bind entities to their attributes Feng and Steinhardt (2023).
+For example, given Sample 1 and 1, a model must bind the entities (e.g., “Box Z”, “Box M”, “Box H”, “Alex”, “John” and “Carl”) to their corresponding attributes (e.g., “coffee”, “stone”, “map”, “bean”, “pie” and “fruit”) so as to recall (or answer) such as what is in “Box Z” or what is sold by “Alex” without confusion.
 Binding has also been studied as a fundamental problem in Psychology Treisman (1996).
-To uncover how Language Models (LMs) realize binding in term of internal representation, Feng and Steinhardt (2023) introduce a Binding ID mechanism and state that LMs apply a abstract concept called Binding ID (BI) to bind and mark Entity-Attribute (EA) pairs (e.g., “Box Z” and “coffee”, as shown in Sample 1, where BI is denoted as a numbered square).
+Figure 1: Our main finding on Ordering ID (OI) subspace intervention.
+Patching entity (e.g., "Z") representations along OI direction (i.e., PC1) in activation space yields corresponding changes in model output.
+To uncover how Language Models (LMs) realize binding in term of internal representation, Feng and Steinhardt (2023) introduce a Binding ID mechanism and state that LMs apply a abstract concept called Binding ID (BI) to bind and mark Entity-Attribute (EA) pairs (e.g., “Box Z” and “coffee” in Sample 1, where BI is denoted as a numbered square).
 They also claim that the BI is represented as a vector to be added on the representation (or activation) of a EA pair so that the common vector is used as a key clue to search attribute for a given entity.
-However, they have not directly captured the prototype of BI information (called BI information hereafter for simplicity) from the activations.
+However, they have not captured the Ordering ID (OI) information from the entity (or attribute) activations that causally affects binding behaviour and thus BI information as well.
+Here, OI is defined as the input order (or ordering index) of entities and attributes, no matter they are bound by a relation (e.g., “is_in” in Sample 1) or not, such as the indexing number in Sample 1 and Sample 1.
+We can observe that in a 1E-to-1A bound context, such as in Sample 1, BI and OI are interchangeable.
 […]
@@
 […]
-Figure 1: Our main finding on Binding ID (BI) subspace intervention. Patching entity (e.g., "Z") representations along BI direction in activation space yields corresponding changes in model output.
 Since binding is the foundational skill that underlies entity tracking Feng and Steinhardt (2023), in this work, we take the entity tracking task Kim and Schuster (2023); Prakash et al.
-(2024) as a benchmark to evaluate the LM’s binding behaviour.
-Based on the analysis of internal representation on this task, we localize the BI information from the activiations and provide a novel view of the Binding ID mechanism.
-Specifically, we discover that LMs encode (or store) BI information into a low-rank subspace (called BI subspace hereafter) that primarily encodes the order (i.e., the input order from left to right) of entity and attributes and which is used as the prototype of BI to determine binding.
-To identify the BI subspace, we take Principle Component Analysis (PCA) as our first attempt  [^1] Besids PCA, we also attempt partial least squares regression for capturing BI subspace.
-Since they achieve similar regression score, we adopt PCA for simplicity.
-See Appendix (§A.2) for details.
-to capture the subspace representing the BI information, and it is empirically proven be effective.
-Therefore, our findings confirm and extend prior understanding of BI in LMs.
-Going beyond the prior work on BI mechanism Feng and Steinhardt (2023), we show that by causally intervening along the BI encoding Principle Component (PC), LMs swap the binding and infer a new attribute for a given entity accordingly.
-That is, we find a consistent causal relationship between the BI subspace intervention and the inferred attributes by LMs.
-For example, as shown in Figue 1, by patching activations along a direction (i.e., PC1), we can make the LMs to infer "Box Z contains the stone" and "Box Z contains the map" instead of "Box Z contains the coffee".
-Overall, our findings suggest that LMs encodes BI information into a subspace of entity activations that primarily encodes the appearance order (or reversed order) of a EA pair in a given context.
-Moreover, the discovered BI subspace plays a crucial role in the in-context binding computation.
-In addition, we find that BI subspace not only exists in the Pretrained large LMs such as Llama2 Touvron et al.
-(2023) and Llama3 AI@Meta (2024), but also in the code fine-tuned LM such Float-7B Prakash et al.
+(2024) as a benchmark to analyze the LM’s binding behaviour.
+Based on the analysis of internal representation on this task, we localize the OI information from the activiations and provide a novel view of the BI mechanism.
+Specifically, we apply Principle Component Analysis (PCA) as well as other dimension reduction methods such as Independent Component Analysis (ICA)  [^1] See Appendix (§A.2) and Appendix (§A.3) for details.
+to analyze the activations of LMs, and which are empirically proven to be effective.
+We discover that LMs encode (or store) the OI information into a low-rank subspace (called OI subspace hereafter), and the discovered OI subspace can causally affect binding behaviour and thus BI information as well.
+That is, we find that by causally intervening along the OI encoding Principle Component (PC), LMs swap the binding and infer a new attribute for a given entity accordingly.
+For example, as shown in Figure 1, by patching activations along the direction (i.e., PC1), we can make the LMs to infer “Box Z contains the stone” and “Box Z contains the map” instead of “Box Z contains the coffee”.
+Therefore, our findings extend the previous BI based understanding of binding in LMs Feng and Steinhardt (2023) by revealing the causality between OI and binding.
+Overall, our findings suggest that LMs encode OI information into a subspace of LMs’ activations that primarily encodes the order index of entities and attributes in a given context.
+What is more, the discovered OI subspace plays a crucial role in the in-context binding computation.
+In addition, we find that such OI subspace that determines binding is prevalent across multiple LM families such as Llama2 Touvron et al.
+(2023) (and Llama3 AI@Meta (2024)), Qwen1.5 Bai et al.
+(2023) and Pythia Biderman et al.
+(2023), and the code fine-tuned LM Float-7B Prakash et al.
 (2024).
-2 Finding Low-rank Subspace for Binding ID
-In this section we describe our Principle Component Analysis (PCA) based method to locate the BI subspace in activations of LMs.
-Firstly we extract entity activation from LMs as shown in Figure 1.
-Given a LM (e.g., Llama2), and a collection of texts which describe a set of EA pairs for a relation such as Sentence 1 for relation “is_in”, we extract the activation of entity token (e.g., “Z”) in query (denoted as $\mathbf{x}_{i}$) from certain layer  [^2] See Appendix (§A.3) for the layer selection.
+2 Finding OI Subspace
+In this section we describe our Principle Component Analysis (PCA) based method to localize the OI subspace in activations of LMs.
+As shown in Figure 1, we firstly extract entity activation from LMs.
+Given a LM (e.g., Llama2), and a collection of texts which describes a set of EA pairs related by a relation such as “is_in” in Sample 1, we extract the activation of entity token (e.g., “Z”) in query (denoted as $\mathbf{x}_{i}$) from certain layer  [^2] See Appendix (§A.4) for the layer selection.
 and construct a activation matrix $M_{r}\in R^{n\times d}$ for a relation $r$, where $n$ denotes the number of entities and $d$ denotes the dimension of the activation.
@@
 (2023).
-Inspired by these studies, we choose PCA as our first attempt to localize BI subspace.
-We hypothesize that in a activation subspace, entities with the same BI tend to cluster together (w.r.t the ones with different BIs), even though these entities have different semantic meaning, and the BIs are encoded as directions (or a PC) in the subspace.
-For convenience, we number BIs in left-to-right order, and the leftmost BI$=0$.
-To capture the subspace, or BI direction, we leverage PCA, which identifies the principle directions of a space.
+Inspired by these studies, we choose PCA as our first attempt to localize OI subspace.
+In addition, we also apply other dimension reduction methods such as Independent Component Analysis (ICA) to capture OI subspace, and which are empirically proven to be effective.
+See Appendix (§A.2) and Appendix (§A.3) for details.
+We hypothesize that in a activation subspace, entities with the same OI tend to cluster together (w.r.t the ones with different OIs), even though these entities have different semantic meaning, and the OIs are encoded as directions (or a PC) in the subspace.
+For convenience, we number OIs in left-to-right order, and the leftmost OI$=0$.
+Since PCA is applied to identify the principle directions from a multidimensional space, we leverage PCA to capture OI subspace (or direction) from activations of LMs.
 Specifically, the PCA of a activation matrix is ${M_{r}}=U_{r}\Sigma_{r}V^{T}_{r}$, where the columns of $V_{r}\in R^{d\times d}$ are principle directions of $M_{r}$.
-We takes first $c$ columns of $V_{r}$ as the BI direction, denoted as $B_{r}\in R^{d\times c}$.
-3 BI Subspace Visualization
-Figure 2: Layer-wise BI subspace visualization on Llama2-7B.
+We takes first $c$ columns of $V_{r}$ as the OI direction, denoted as $B_{r}\in R^{d\times c}$.
+3 OI Subspace Visualization
+Figure 2: Layer-wise OI subspace visualization on Llama2-7B, where “BI” primarily denotes OI.
 We adopt a subset of the entity tracking dataset Kim and Schuster (2023); Prakash et al.
-(2024), which contains $n=1000$ samples, to create layer-wise activation matrix $M_{r}^{l}$.
-We then uses the $M_{r}^{l}$ to extract the layer-wise BI subspace projection matrix $B_{r}^{l}\in R^{d\times 2}$ to visualize the activation.
-Figure 2 shows the embedding visualization on Llama2-7B, where each point represents the activation of an entity projected via the $B_{r}^{l}$, and the colors represent BIs.
-From which, we can observe that middle layers, such as layer 8, have a clearly visible direction along which BI increases (or decreases), while the others have tangled distribution.
-We also observe similar property of distribution on Llama3-8B and Float-7B (§A.4).
-This indicates that LMs use the middle layers to encode BI information.
-We call this dimension that represents the order of BI as BI Principle Component (BI-PC).
-In the following section, we apply causal intervention on the BI-PC to analyze how BI-PC affect the model output.
-4 Causal Interventions on Binding ID Subspace
+(2024), which contains $n=1000$ samples, to create layer ($l$) wise activation matrix $M_{r}^{l}$.
+We then uses the $M_{r}^{l}$ to extract the layer-wise OI subspace projection matrix $B_{r}^{l}\in R^{d\times 2}$ to visualize the activations.
+Figure 2 shows the embedding visualization on Llama2-7B, where each point represents the activation of an entity projected via the $B_{r}^{l}$, and the colors represent OIs.
+From which, we can observe that middle layers, such as layer 8, have a clearly visible direction along which OI increases, while the others have tangled distribution.
+We also observe similar pattern of distribution on Llama3-8B, Float-7B (§A.5) and other LM families such as Qwen1.5 and Pythia (§A.6).
+This indicates that LMs use the middle layers to encode OI information, and the finding is prevalent across multiple LM families.
+This finding is also consistent with the “stages of inference hypothesis” Lad et al.
+(2024) stating that the function of early layers is to perform detokenization, middle layers do feature engineering, and late layers map the representations from the middle layers into the output embedding space for next-token prediction.
+According to the hypothesis, we would expect to find the ordering feature most prominently represented in middle layers, which is exactly what the visualization shows.
+We call this dimension that represents OI as OI Principle Component (OI-PC).
+In the following section, we apply causal intervention on the OI-PC to analyze how OI-PC affect the model output.
+4 Causal Interventions on OI Subspace
 Answer for # Step
@@
 Box Z contains the ice ring ring clock paper engine
-Table 1: Attributes inferred by Llama2-7B as a result of directed activation patching along BI-PC in the BI subspace on the dataset of “r: is_in”, where color denotes the BI.
-By projecting the activation matrix $M_{r}$ into the BI subspace, we have found a correlative evidence for the existence of the direction (i.e., BI-PC) that encodes Binding IDs.
-However, it is possible that the BI information is encoded in the BI subspace but has no effect on model output.
-In order to test if Binding IDs are not only encoded in the BI subspace, but that these representations can be steered so as to swap the binding and change LM’s output.
-We now perform interventions to establish the causality.
-That is, we want to find out if making BI swapping interventions leads to a change in model output.
+Table 1: Attributes inferred by Llama2-7B as a result of directed activation patching along OI-PC in the OI subspace on the dataset of “r: is_in”, where color denotes the BI.
+By projecting the activation matrix $M_{r}$ into the OI subspace, we have found a correlative evidence for the existence of the direction (i.e., OI-PC) that encodes OI information.
+However, it is possible that the OI information is encoded in the OI subspace but has no effect on LMs’ binding behaviour.
+In order to test if OIs are not only encoded in the OI subspace, but that these representations can be steered so as to swap the binding and change LM’s output, in this section, we perform interventions to analyze the causality.
+That is, we want to find out if making interventions along OI-PC leads to a change in LM’s binding computation.
 4.1 Activation Patching
-Since LM computation graph could be viewed as causal graph Meng et al.
-(2022); McGrath et al.
-(2023), we intervene on model activations via activation patching Vig et al.
-(2020); Wang et al.
-(2022); Zhang and Nanda (2023); Heinzerling and Inui (2024); Engels et al.
-(2024), and observe the effect on model output.
-Unlike the common activation patching setup in which one replaces activations resulting from an original input with activations from a corrupted one, we create patches by editing activations along a particular direction (i.e., along BI-PC), similar to the activation editing method of Matsumoto et al.
+Activation Patching (AP) Vig et al.
+(2020); Geiger et al.
+(2020, 2021); Wang et al.
+(2022); Stolfo et al.
+(2023); Heinzerling and Inui (2024); Engels et al.
+(2024); Hanna et al.
+(2024) has been recently proposed to causally intervene computational graph of a LM so as to interpret the function of a target computational node (or edge).
+The process of AP usually involves preparing corrupted input, using its activations to replace the corresponding ones obtained from original input and analyzing the effect on model output.
+Different with the common AP setup, we realize AP by directly editing activations along a particular direction (i.e., along OI-PC), similar to the activation editing method of Matsumoto et al.
 (2023); Heinzerling and Inui (2024); Engels et al.
 (2024).
-Although automatic methods for localizing model circuits of interest have been proposed Conmy et al.
-(2023); Kramár et al.
-(2024), for simplicity we perform a coarse layer-wise search based on the effect of activation patching in a development set, as shown in Appendix (§A.3), and use the found setting for all experiments.
 4.2 Setting
@@
 (2024).
-The dataset contains English sentence describing a set of objects located in a set of boxes with difference labels, and the task is to infer what is inside a given box.
-For instance, when a LM is presented with “The coffee is in Box Z, the stone is in Box M, the map is in Box H, … Box Z contains the”, it should infer the next token as “coffee”.
-Each sample involves $7$ AE pairs.
+The dataset consists of English sentence describing a set of objects (here called attributes) located in a set of boxes with difference labels (here called entities), and the task is to infer what is contained by a given box.
+For instance, when a LM is presented with “The coffee is in Box Z, the stone is in Box M, the map is in Box H, … Box Z contains the”, the LM should infer the next token as “coffee”.
+Each sample involves $7$ EA pairs.
 To evaluate the binding in various context, we also apply the templates shown in Table 2 to generate other $5$ datasets with different relation, where $a_{i}$ and $e_{i}$ denotes the attribute and entity, and they are sampled from a fixed pool of $224$ one-token objects (e.g., “dog”, “corn” and “cookie”) and $523$ of one-token names (e.g., “Alex”, “Juli” and “Dan”) respectively.
@@
 Table 2: Templates of Dataset.
-Metrics We apply two evaluation metrics.
-The logit difference metric introduced in Wang et al.
-(2022), which calculates difference in logits between original and intervened answers, as well as the "logit flip" accuracy metric Geiger et al.
-(2022), which represents the proportion of cases where we alter the model output after a causal intervention.
+Metrics We apply two evaluation metrics: logit difference and logit flip.
+The logit difference metric is introduced in Wang et al.
+(2022), which calculates difference in logits of a target token between original and intervened setting.
+The "logit flip" accuracy metric is introduced by Geiger et al.
+(2022), which represents the proportion of candidate tokens in model output after a causal intervention.
 (a) r: is_in
@@
 (f) r: push
-Figure 3: Logit Difference (LD) for BI-PC based intervention across datasets on Llama2-7B, where x axis denotes the number of intervention steps on $e_{0}$, y axis does the LD, BI_i represents each target attribute and the light yellow bottom line indicates the LD of original attribute (i.e., $a_{0}$).
+Figure 3: Logit Difference (LD) for OI-PC based intervention across datasets on Llama2-7B, where x axis denotes the number of intervention steps on $e_{0}$, y axis does the LD, BI_i represents each target attribute and the light yellow bottom line indicates the LD of original attribute (i.e., $a_{0}$).
 Here, $l=8$, $v=2.5$, and $\alpha=3.0$.
-4.3 Results: Direct Editing BI Subspace
-We hypothesize that LMs encode BI information into a low-rank BI subspace.
-Therefore, we wonder if a LM changes the binding behavior, when adding a particular value $v$ (called step hereafter) along the BI-PC mentioned in Section (§2).
-For example, if we add one unit of $v$ on the activation of $e_{0}$, the LM will reset its BI as $1$ and bind attribute $a_{1}$ to entity so that infers the $a_{1}$ as the attribute of $e_{0}$ instead of the original $a_{0}$.
+4.3 Results: Direct Editing OI Subspace
+We hypothesize that LMs encode OI information into a low-rank subspace that causally affect BI information and thus binding behaviour as well.
+Therefore, we wonder if a LM changes the binding behavior, when adding a particular value $v$ (called step hereafter) along the OI-PC mentioned in Section (§2333where PC2 is assigned a fixed value, which is a hyperparameter..
+For example, if we add one unit of $v$ on the OI-PC of $e_{0}$, the LM will reset its BI as $1$ and bind attribute $a_{1}$ to the entity based on the similarity of BI so that infers the $a_{1}$ as the attribute of $e_{0}$ instead of the original $a_{0}$.
 Similarly, adding two units of $v$ will make the LM infer $a_{2}$ for $e_{0}$, and so on.
-We intervene via the Equation 1, where $\mathbf{x}_{0,l}$ is the original activation of $e_{0}$ (i.e., the leftmost entity) in layer $l$, $\mathbf{x}_{0,l}^{*}$ is the intervened activation, $B_{r}$ is the BI subspace projection matrix mentioned in Section (§2), $\alpha$ is a hyper-parameter to scale the effect of intervention and $\beta$ ($0\leq\beta\leq 6$) denotes the number of steps.
+We intervene via the Equation 1, where $\mathbf{x}_{0,l}$ is the original activation of $e_{0}$ (i.e., the leftmost entity) in layer $l$, $\mathbf{x}_{0,l}^{*}$ is the intervened activation, $B_{r}$ is the OI subspace projection matrix mentioned in Section (§2), $\alpha$ is a hyper-parameter to scale the effect of intervention and $\beta$ ($0\leq\beta\leq 6$) denotes the number of steps.
 $\mathbf{x}_{0,l}^{*}=\mathbf{x}_{0,l}+\alpha B_{r}^{T}(B_{r}\mathbf{x}_{0,l}+\beta v)$ (1)
-Table 1 lists several examples under the BI subspace intervention on the entity tracking dataset Kim and Schuster (2023); Prakash et al.
+Table 1 lists several examples under the OI subspace intervention on the entity tracking dataset Kim and Schuster (2023); Prakash et al.
 (2024).
 We also list the examples from other datasets in Appendix (§A.7).
-We can see that when adding $1$ step along BI-PC, the model selects “stone” for entity “Z” instead of its original attribute “coffee”.
+We can see that when adding $1$ step along OI-PC, the model selects “stone” for entity “Z” instead of its original attribute “coffee”.
 Similarly, when the step is doubled, the model will select attribute “map” for the entity, and so on.
-Although the attribute selection does not strictly follow the number of steps, this indicates, to some extent, that changing the value along BI-PC can induce the swap of attribute.
-Besides the qualitative analysis, we also conduct quantitative analysis for the causality between the BI subspace based activation patching and the binding behaviour of the LM.
-We plot mean-aggregated effect of directed activation patching across multiple datasets in Figure 3.
+This indicates that changing the value along OI-PC can induce the swap of attribute.
+In addition, we notice that some attributes are repeated or skipped after the intervention.
+The reason is that, as shown in Figure 2, OI is represented as a continuous range (e.g., $[a_{0},b_{0}],[a_{1},b_{1}],[a_{2},b_{2}],...$) on OI-PC, implying the points in the same range share OI, and if a sample (e.g., its OI-PC is $s_{i}$) is still in the range (e.g., $a_{0}<(s_{i}+v)<b_{0}$) or skip its neighbouring range (e.g., $a_{2}<(s_{i}+v)<b_{2}$) after the intervention (e.g., $s_{i}+v$), its OI will be the same or added by $2$, and thus the binding information will change accordingly.
+Besides the qualitative analysis, we also conduct quantitative analysis for the causality between the OI subspace based AP and the binding behaviour of LMs.
+We plot mean-aggregated effect of the OI-PC based AP across multiple datasets in Figure 3.
 Figure 3 indicates how the Logit Difference (LD) of each attribute changes as the step increases.
@@
 In contrast, LD of other attributes gradually increase until a certain point and then gradually decrease.
-Given a candidate attribute, its peak point generally corresponds the number of steps that equals to its BI.
-For instance, when adding $3$ steps, the points of BI_3 (i.e., attributes of BI$=3$) on step$=3$ achieve the highest LD score.
-This indicates that by adjusting the value along the BI-PC, we can increase the probability of the corresponding attribute, thereby swap the answer.
+Given a candidate attribute, its LD peak roughly corresponds to the number of steps that is equal to its BI.
+For instance, when adding $3$ steps, the points of BI_3 (i.e., attributes of BI$=3$) achieve the highest LD score.
+This indicates that by adjusting the value along the OI-PC, we can adjust BI information and thus increase the logit score of the corresponding attribute.
 Similarly, Figure 4 illustrates the relation between the number of steps and the logit flip, which gauges the percentage of the predicted attributes under an intervention.
-Figure 4 shows that as the step increases, the proportion bar becomes darker, it means that the model promotes the proportion of the corresponding following attribute in its inference.
+Figure 4 shows that as the step increases, the proportion bar becomes darker, it means that the model promotes the proportion of the corresponding attribute in its inference.
 For instance, when adding $3$ step on the subspace, the $a_{3}$ (i.e., BI_3) becomes the major of the answers.
-This proves that the discovered subspace stores BI information, and the subspace will causally affect the computation of Binding in a LM.
-See Appendix (§A.9) for the results on Llama3-8B.
+This proves that the OI-PC based interventions can causally affect BI information as well as the computation of Binding in a LM.
+See Appendix (§A.10) for the results on Llama3-8B, Appendix (§A.6) for the results on Qwen1.5-7B and Pythia-6.9B, and Appendix (§A.9) for the results on Float-7B.
 (a) r: is_in
@@
 (f) r: push
-Figure 4: Logit flip for BI-PC based intervention across datasets on Llama2-7B, where x axis denotes the number of intervention steps on $e_{0}$, y axis does the proportion of each inferred attribute in model output.
-4.4 Results: Activation Steering on BI Subspace
+Figure 4: Logit flip for OI-PC based intervention across datasets on Llama2-7B, where x axis denotes the number of intervention steps on $e_{0}$, y axis does the proportion of each inferred attribute in model output.
+4.4 Results: Activation Steering on OI Subspace
 Figure 5: Logit Difference and Logit Flip for activation steering on the entity tracking dataset (i.e., r: is_in), where x axis represents the intervention of $\mathbf{s}_{0\rightarrow bi}$.
 Inspired by the research on Activation Steering (AS) Turner et al.
-(2023), we apply an AS method to verify the importance of the BI subspace on LM’s binding behaviour.
+(2023), we apply an AS method to verify the importance of the OI subspace on LM’s binding behaviour.
 Specifically, we use the following Equation 2 to extract a subspace steering vector $\mathbf{s}_{0\rightarrow bi}$, which is proposed to swap BI from 00 to $bi$, where $n$ is the number of target entities, $\mathbf{x}_{bi,l}^{i}$ represents the activation of entity $e_{i}$ from layer $l$, and its BI is $bi$.
-We intervene via Equation 3 and assume that by adding $\mathbf{s}_{0\rightarrow bi}$ to the original activation $\mathbf{x}_{0,l}$, we can increase the LD and the proportion of the attribute $a_{bi}$.
+We intervene via Equation 3 and assume that by adding $\mathbf{s}_{0\rightarrow bi}$ on the original activation $\mathbf{x}_{0,l}$, we can increase the LD and the proportion of the attribute $a_{bi}$.
 Figure 5 shows the results on the entity tracking dataset Kim and Schuster (2023); Prakash et al.
@@
 (Appendix (§A.8) shows the results on other datasets) These results indicate that AS can achieve the similar tendency as the direct value intervention mentioned in Section (§4.3).
-For instance, adding $\mathbf{s}_{0\rightarrow 3}$, which is used to swap BI form 00 to $3$, can increase the LD of $a_{3}$ and its proportion in the predicted answers.
-The consistent tendency with the results of the direct subspace editing, shown Figure 3 and Figure 4, further illustrates that the discovered subspace contains the BI information, and more importantly, it plays an important role when the model performs in-context binding computation.
+For instance, adding $\mathbf{s}_{0\rightarrow 3}$, which is supposed to swap BI from 00 to $3$, can increase the LD of $a_{3}$ and its proportion in the predicted answers.
+The consistent tendency with the results of the direct subspace editing, shown Figure 3 and Figure 4, further illustrates that the discovered OI subspace can causally affect BI information and binding behaviour.
+Therefore, OI subspace plays an important role when LMs perform in-context binding computation.
 $\mathbf{s}_{0\rightarrow bi}=\frac{1}{n}\sum_{i=1}^{n}(B_{r}\mathbf{x}_{bi,l}^{i}-B_{r}\mathbf{x}_{0,l}^{i})$ (2)
 $\mathbf{x}_{0,l}^{*}=\mathbf{x}_{0,l}+\alpha B_{r}^{T}\mathbf{s}_{0\rightarrow bi}$ (3)
-4.5 Binding Subspace and Position
+4.5 OI Subspace and Position
 Input (original)
@@
 $C^{\prime}_{2}$ $a_{*0}^{p0}\,r\,e_{*0}^{p1},\,a_{1}^{p2}\,r\,e_{1}^{p3},\,a_{2}^{p4}\,r\,e_{2}^{p5}.\quad e_{2}^{p5}\ r^{-1}\ ?$
-Table 3: Simplified expression of original inputs and the one modified with pseudo relation, which is proposed to equalize PI for PCA analysis, where $a_{0}^{p1}\,r\,e_{0}^{p2}$ represents a relation such as “the apple is in Box C”, and $e_{0}^{p2}$ denotes an entity with BI of 00 and PI of $p2$, $e_{2}^{p5}\ r^{-1}\ ?$ denotes the query on entity $e_{1}$, such as “Box C contains the”.
-Input (original)
-$a_{0}^{p0}\,r\,e_{0}^{p1},\,a_{1}^{p2}\,r\,e_{1}^{p3},\,a_{2}^{p4}\,r\,e_{2}^{p5}.\quad e_{0}^{p1}\ r^{-1}\ ?$
-Input (with interjection)
-$a_{0}^{p0}\,r\,e_{0}^{p1},\,j^{p2}\,j\,j^{p3}\,...j^{pi}\,a_{1}^{pi+1}\,r\,e_{1}^{pi+1}...\quad e_{0}^{p1}\ r^{-1}\ ?$
-Table 4: Simplified expression of original input and the one modified with a sequence of interjections, where $j^{p3}$ denotes an interjection $j$, such as “ah”, with position of $p3$, which is also the position of $e_{1}^{p3}$ in the original input.
+Table 3: Simplified expression of original inputs and the one modified with pseudo relation, which is proposed to equalize PI for PCA analysis, where $a_{0}^{p1}\,r\,e_{0}^{p2}$ represents a relation such as “the apple is in Box C”, and $e_{0}^{p2}$ denotes an entity with OI of 00 and PI of $p2$, $e_{2}^{p5}\ r^{-1}\ ?$ denotes the query on entity $e_{1}$, such as “Box C contains the”.
 Input (original)
@@
 Table 4: An example of the dataset with filler words “I will find out that”.
-In this section, we discuss the relationship between the BI subspace and Positional Information (PI), namely the $postion\_ids$ of input tokens.
-As mentioned in Section (§4.3), the discovered subspace stores the BI information, therefore direct intervention on it can swap the answer of LM.
-However, one counter hypothesis is that the subspace is not used for storing BI information but the PI of attributes, thus the direct intervention merely change the PI so that swap the answer.
-Regarding the relationship between BI and PI, Feng and Steinhardt (2023) found that even when PI of attributes is swapped, the model still makes correct predictions, thus confirming the independence between BI and PI.
-Based on this finding, we go one step further and illustrate the independence between the BI subspace and PI.
-To prove the independence, we create three datasets, the first one is by extending the original dataset with pseudo relation, as shown in Table 3, the second one is by inserting a sequence of interjection(e.g., “ah, ah, …”), as listed in Table 4, and the third one is by prefixing the original dataset with filler words(e.g., “It can be seen that”).
+In this section, we discuss the relationship between the OI subspace and Positional Information (PI), which is namely the $postion\_ids$ of input tokens.
+As mentioned in Section (§4.3), the discovered subspace can causally determine BI information.
+Therefore, direct intervention on the subspace can swap the answer of a LM.
+However, one counter hypothesis is that the subspace is not used for storing OI information but the PI of attributes, and thus the direct intervention merely changes the PI of answer token so that the LM applies the new PI to generate corresponding token.
+In other words, the OI space (or OI-PC) might have high correlation with PI but not OI.
+To prove the independence between OI subspace and PI, we create three datasets, the first one is by extending the original dataset with pseudo relation, as shown in Table 3, the second one is by prefixing the original dataset with filler words(e.g., “It can be seen that”) and third one is by inserting a sequence of interjection(e.g., “ah, ah, …”), as listed in Table 12.
 New Dataset with Pseudo Relation.
-In Table 3, $a_{*0}^{p0}\,r\,e_{*0}^{p1}$ refers to a pseudo relation, which is a fixed expression, such as “the PC is in Box Z”, applied to adjust the PI while keeping the BI.
-For instance, in Table 3, adding one or two $a_{*0}^{p0}\,r\,e_{*0}^{p1}$ before $a_{1}\,r\,e_{1}$ (i.e., $C^{\prime}_{2}$ and $C^{\prime}_{1}$) does not affect the BI of $e_{1}$ but its PI, because $e_{1}$ is still the second unique entity from the left (i.e., BI$=1$), but its PI is $p3$ and $p5$ respectively.
-Using the pseudo relation, we create the data in a manner that the target entity (e.g., $e_{1}^{p5}$ and $e_{2}^{p5}$) to extract activation have the same PI, such as $C^{\prime}_{1}$ and $C^{\prime}_{2}$ in Table 3.
-We apply the method mentioned in Section (§2) on the set of activations $\{\overrightarrow{e_{1}^{p5}},\overrightarrow{e_{2}^{p5}},...\}$, where $\overrightarrow{e_{1}^{p5}}$ denotes the activation of $e_{1}$ in $e_{1}^{p5}\ r^{-1}\ ?$, so as to capture the BI difference and exclude the PI difference, because they share the same PI (i.e., $P5$) but different BI (i.e., $1$, $2$, …).
-Then we compare its BI subspace with the original one (e.g., $\{\overrightarrow{e_{1}^{p2}},\overrightarrow{e_{2}^{p5}},...\}$) to analyze how the distribution of BI subspace changes after removing the PI variance.
-Figure 6 visualizes the BI subspace distribution, where the light colored points denote the original distribution, and the dark ones are from the new one with equalized PI.
-We can observe that after removing the PI difference, the distribution is still similar to the original one that there is a clearly visible direction along which BI increases.
-This illustrates that our PCA based method can capture BI information, that is, along the direction of BI-PC, and it does not causally depend on PI.
-See Appendix (§A.5) for our further analysis on how the context of binding affects the distribution.
-New Dataset with Interjections.
-The dataset is created by inserting a sequence of interjections after the first attribute entity pair, as illustrated in Table 4.
-Since there is no BI information in the interjection (e.g., $j^{p3}$), adding it only changes the PI of its following entities and attributes.
-We set the number of interjections as that its last PI is larger than the last PI of its original input (e.g., $pi>p5$ in Table 4).
-Based on this dataset, we conduct the same intervention on its BI subspace, as mentioned in Section (§4.3).
-The counter argument is that the subspace only captures PI, and the intervening step only changes the PI information.
-Specifically, adding one unit of $v$ on $\overrightarrow{e_{0}^{p1}}$ can change its PI from $p1$ to $p3$, which is the PI of $e_{1}$, and its attribute is $a_{1}$, as shown in Table 4, thereby the LM swaps the answer from $a_{0}$ to $a_{1}$.
-If it is true, then the same intervention will not change the answer on the new dataset, because following the counter argument, after adding one unit of $v$ on $\overrightarrow{e_{0}^{p1}}$, its PI becomes $p3$, and $p3$ is the PI of $j^{p3}$.
-The LM thus would not select $a_{1}$ as its answer.
-However, the results on Figure 7 shows that the subspace intervention on the new dataset achieves similar results as the original one, as shown in Figure 3 and Figure 4, proving the counter argument wrong and indicating the independence between BI subspace and PI.
+In Table 3, $a_{*0}^{p0}\,r\,e_{*0}^{p1}$ refers to a pseudo relation, which is a fixed expression, such as “the PC is in Box Z”.
+The pseudo relation is applied to adjust the PI while keeping the OI.
+For instance, in Table 3, adding one or two $a_{*0}^{p0}\,r\,e_{*0}^{p1}$ before $a_{1}\,r\,e_{1}$ (i.e., $C^{\prime}_{2}$ and $C^{\prime}_{1}$) does not affect the OI of $e_{1}$ but its PI, because $e_{1}$ is still the second unique entity from the left (i.e., its ordering index 0I$=1$), but its PI is $p3$ and $p5$ respectively.
+Using the pseudo relation, we create the data in a manner that the target entity for activation analysis (e.g., $e_{1}^{p5}$ and $e_{2}^{p5}$) have the same PI but different OI, such as $C^{\prime}_{1}$ and $C^{\prime}_{2}$ in Table 3.
+We apply the method mentioned in Section (§2) on the set of activations $\{\overrightarrow{e_{1}^{p5}},\overrightarrow{e_{2}^{p5}},...\}$, where $\overrightarrow{e_{1}^{p5}}$ denotes the activation of $e_{1}$ in query (i.e., $e_{1}^{p5}\ r^{-1}\ ?$), so as to capture the OI difference and exclude the PI difference, because they share the same PI (i.e., $P5$) but different OI (i.e., $1$, $2$, …).
+Then we compare its OI subspace with the original one (e.g., $\{\overrightarrow{e_{1}^{p2}},\overrightarrow{e_{2}^{p5}},...\}$) to analyze how the distribution of OI subspace changes after removing the PI variance.
+Figure 6 visualizes the OI subspace distribution, where the light colored points denote the original distribution, and the dark ones are from the new one with equalized PI.
+We can observe that after removing the PI difference, the distribution is still similar to the original one that there is a clearly visible direction along which OI increases.
+This illustrates that our PCA based method can capture OI information, that is, along the direction of OI-PC, and it does not causally depend on PI.
 New Dataset with Filler Words.
@@
 (2024), as shown in Table 4.
-Since the length (i.e., the number of tokens) of FW directly affects the PI of its following entities and attributes without affecting their BI, we take the length as the measure of intervention on PI and apply Spearman’s rank correlation $\rho$ to calculate the correlation between the length (denoted as PI) and the discovered BI-PC.
-Figure 8 shows $\rho$ between PI and BI-PC as well as between BI and BI-PC.
-We can observe that BI-PC has high $\rho$ with BI but almost zero $\rho$ with PI, indicating that the discovered BI-PC is dependent on BI but independent on PI.
-Therefore, the BI-PC does not simply encode absolute token position.
+Since the length (i.e., the number of tokens) of FW directly changes the PI of its following entities and attributes without affecting their OIs, we take the length as the measure of intervention on PI and apply Spearman’s rank correlation $\rho$ to calculate the correlation between the length (denoted as PI) and the OI-PC value.
+Figure 7 shows $\rho$ between PI and OI-PC as well as between OI and OI-PC.
+We can observe that OI-PC has high $\rho$ with OI but almost zero $\rho$ with PI, indicating that the discovered OI-PC is highly correlated with OI information but independent on PI.
+Therefore, the OI-PC does not simply encode absolute token position.
+See Appendix (§A.11) for the analysis on the New Dataset with Interjections.
 Figure 6: Embedding visualization for activation with equalized PI, where “Ori” denotes the distribution of original dataset, while “Pse” denotes the distribution of the new dataset with pseudo relation.
-Figure 7: Logit Difference and Logit Flip for activation patching on the dataset of “r: is_in” with interjections.
-Appendix (§A.10) shows the results on other dataset with the same interjection based modification.
 […]
-4.6 Binding Subspace and Relational Information
+4.6 Consistency of OI Subspace
+The BI mechanism Feng and Steinhardt (2023) mentions that a LM represents a related EA pair through the consistency of their BI information.
+This naturally raises a research question that if there is the consistency between OI-PC of entities (e.g., OI-PC of “Box Z” in Sample 1) and of their corresponding attributes  [^4] The OI-PC of attribute is extracted from the query of attribute such as “(context) The coffee is in”.
+(e.g., OI-PC of“coffee”).
+To analyze the consistency of OI-PC between a related EA pair, we prepare four alternative datasets with various binding pattern as shown in Table 6.
+We use the distance on OI-PC to measure consistency and show partial result of mutual OI-PC based distance in Figure 8.
+We can observe that potentially related EA pairs (e.g., “E0_0” and “A0_0”) tend to have lower OI-PC based distance than arbitrary pairs in all binding patterns, indicating that to some extent, OI-PC based distance could be seen as an important feature to represent binding.
+To further illustrate its significance, we attempt to classify related EA pairs only relying on their OI-PC based distances.
+Specifically, we search an optimal threshold value from a development set, and which is applied to classify potentially related EA pairs in a testing set.
+The results are shown in Figure 9.
+We can observe that the performance of the OI-PC (i.e., “PC1”) based distance is significantly better than other PCs across all binding patterns, indicating that comparing to other PCs, the OI-PC could be used to compute binding information.
+4.7 OI Subspace and Relatedness
 Input (original)
@@
 Input (Non-related)
-“I see apple, cup and hair, somewhere else there is Box E, the bell, bus, toy, cookie and Box F are scattered around, ”
+“I see apple, somewhere else there is Box E, the bell and Box F are scattered around, …”
 Table 5: An example of the dataset with non-related expression.
-In order to study the relationship between BI-PC and Relational Information(RI), we create an alternative dataset by converting relational expression of the entity tracking dataset Kim and Schuster (2023); Prakash et al.
+Input (7A-7E)
+“A0 is in E0, A1 is in E1, A2 is in E2, A3 is in E3, A4 is in E4, A5 is in E5, A6 is in E6.”
+Input (7A-3E)
+“A0 is in E0, A1 is in E0, A2 is in E0, A3 is in E1, A4 is in E1, A5 is in E2, A6 is in E2.”
+Input (7A-2E)
+“A0 is in E0, A1 is in E0, A2 is in E0, A3 is in E1, A4 is in E1, A5 is in E1, A6 is in E1.”
+Input (7A-5E)
+“A0 is in E0, A1 is in E0, A2 is in E0, A3 is in E1, A4 is in E2, A5 is in E3, A6 is in E4.”
+Table 6: Simplified expression of the datasets with various binding pattern, where “7A-3E” represents the pattern containing 7 attributes and 3 entities.
+(a) 7A-7E
+(b) 7A-3E
+(c) 7A-2E
+(d) 7A-5E
+[…]
+Binding of a EA pair means that the EA pair is bound by a binding relation such as property and location Treisman (1996).
+In turn, there is no binding when the EA pair is unrelated in its context.
+This raises a research question that if the correlation between OI-PC of attributes (e.g., OI-PC of “apple”) and of their corresponding entities (e.g., OI-PC of “Box E”) represents the relatedness namely the existence of a relation.
+In order to uncover the relationship between OI-PC and the relatedness, we create an alternative dataset by converting relational expression of the entity tracking dataset Kim and Schuster (2023); Prakash et al.
 (2024) into non-related one.
 Specifically, we prepare a set of non-related expression templates and randomly select one to replace the original expression of relation as shown in Table 5.
-We can observe that the template could make a target EA pair (e.g., “apple” and “Box E”) semantically unrelated, but retain their Ordering Information(OI) (i.e., “apple”->“bell”, and “Box E”->“Box F”).
-The BI mechanism Feng and Steinhardt (2023) notes that a LLM represents binding through the consistency of BI.
-According to the mechanism, we adopt the correlation score between BI-PC of attributes (e.g., BI-PC of “apple”) and of their corresponding entities (e.g., BI-PC of“Box E”) to represent their binding strength.
-We select Spearman’s rank correlation $\rho$ as the correlation score and present the $\rho$ of the non-related dataset as well as the related one in the Figure 9.
-We can observe that there is strong correlation between BI-PC of attributes and of their corresponding entities in the original data, indicating that LLM applies the correlation of BI-PC to encode binding.
-We can also observe that $\rho$ of non-related dataset is lower than the related (i.e., original) one, indicating that the BI-PC contains a certain degree of RI so that removing RI can decrease the $\rho$.
-However, even there is no RI but OI in the non-related dataset, the BI-PC of attributes still has high correlation with their co-occurring entities, indicating that BI-PC not only stores RI, but also encodes OI, and the latter seems to be the main and fundamental factor that is utilized by the LLM to determine binding, because intervention on BI-PC can causally change the behaviour of LLM (as discussed in Section (§4).
-Therefore, we hypothesize that the BI-PC primarily encodes the OI of entities and attributes that is used as a prototype of BI to causally determine binding.
+We can observe that the template could make a target EA pair (e.g., “Box E” and “apple”) semantically unrelated but retain their OI (e.g., the OI of “apple” and “bell” are still 00 and $1$ respectively).
+We select Spearman’s rank correlation $\rho$ as the correlation metric and compare the $\rho$ of the non-related dataset with the related one in the Figure 10.
+We can observe that $\rho$ of non-related dataset is slightly lower than the related (i.e., original) one, indicating that the OI-PC might contain limited relational information so that removing it can marginally decrease the $\rho$.
+However, there is still strong correlation between the non-related (or non-bound) entity attribute pair, indicating that the OI-PC primarily encodes the OI information but not binding information specifically the information of binding relation.
 […]
-4.7 Results on Fine-Tuned LM
-Kramár et al.
-(2024); Kim et al.
-(2024) claim that the code fine-tuned LM, such as Float-7B Prakash et al.
-(2024) outperforms the pretrained LM on the entity tracking task Kim and Schuster (2023); Prakash et al.
-(2024).
-Since the code fine-tuned LM performs well on the entity tracking task that requires the BI subspace based computation, we hypothesize that BI subspace also exists in the code fine-tuned LM and the intervention along BI-PC will causally affect the model output.
-To prove the hypothesis, we conduct the intervention on Float-7B and show results in Figure 10.
-We found that the BI subspace based intervention on Float-7B achieves the similar results as on Llama2-7B, indicating that the BI subspace not only exists in the pretrained LM but also in the fine-tuned one.
-In addition, adding the same step value (i.e., $v$) on Float-7B will achieve higher LD value than Llama2-7B, indicating that the code fine-tuned LM is more sensitive to the BI subspace based intervention.
-For instance, the maximum LD of $a_{4}$ in the former is around $10$, and it is $2$ times larger than the one in the latter, which is around $5$.
-This might partially explains why the code fine-tuned LM performs better than the original one, because code fine-tuning might enhance the function of BI subspace so that it is more sensitive on the intervention and more effective on the in-context entity tracking task.
-Figure 10: Logit Difference and Logit Flip for activation patching on the entity tracking dataset (i.e., r: is_in) on Float-7B. See Appendix (§A.8) for the results on other datasets.
+[…]
 5 Related Work
@@
 (2023), and numeric values such as elevation, population, birth year, and death year Gurnee and Tegmark (2023); Heinzerling and Inui (2024).
-Continuing this line of research, in this work, we discover that LMs such as Llama-2 can also linearly encode BI, because there is a linear direction that primarily encodes BI in the activations.
-See Appendix (§A.1) for other related work.
+Continuing this line of research, in this work, we discover that multiple LMs such as Llama2 can also linearly encode OI along a OI increasing direction in the activations.
+See Appendix (§A.1) for additional related work.
 6 Conclusion and Future Work
@@
 Our future work includes: 1.
-the analysis of BI subspace in a more realistic setting; 2.
+the analysis of OI subspace in a more realistic setting; 2.
 the study of interaction between in-context binding and factual knowledge learned from pretraining; 3.
-BI subspace based mechanistic analysis.
-Limitation
-The limitations of our research include the following points.
-Firstly, we only analyze BI subspace on the attribute prediction task, but not on the entity inference task (i.e., given an attribute to infer its entity).
-Secondly, we lack the analysis on how predicate (or relation) affect the BI subspace, and how the results of BI-subspace based intervention differ with the type of predicate.
-Thirdly, although we use a publicly available entity tracking dataset, it is still a synthesized dataset.
-Therefore, for uncovering how LMs bind and track entity in reality, it is necessary to analyze the BI subspace on a real world dataset.
-The last but not lest, we only analyze binding from the perspective of representation and localize BI subspace.
+OI subspace based mechanistic analysis.
+Acknowledgement
+This work was supported by JST CREST Grant Number JPMJCR20D2 and JSPS KAKENHI Grant Number 21K17814. We are grateful to the anonymous reviewers for their constructive comments.
+Limitations
+The limitations of our research include the following points: 1.
+We only analyze OI subspace on the attribute prediction task, but not on the entity inference task (i.e., given an attribute to infer its entity); 2.
+Although PCA based analysis is empirically proven to be effective, we lack the theoretical analysis on why PCA could capture OI subspace; 3.
+We lack the analysis on how predicate (or relation) affect the OI subspace, and how the results of OI subspace based intervention differ with the type of predicate; 4.
+Although we use a publicly available entity tracking dataset, it is still a synthesized dataset.
+Therefore, for uncovering how LMs bind and track entity in reality, it is necessary to analyze the binding via a real world dataset; 5.
+We only analyze the activations in the query part instead of in the context part, thus this work can not explain how LMs encode OI in the context and how it is used for OI encoding in query part and how it contributes the binding computation; 6.
+We infer the change of BI information via the result of model output but the interaction between OI and BI information is not directly observed.
+This research thus lacks the mechanistic interpretation on the interaction between them; 7.
+We only analyze binding from the perspective of representation and localize OI subspace.
 However, we have not answered what is the mechanism that generates the subspace and what is the circuit that utilizes the subspace for binding.
@@
 The existing dataset Kim and Schuster (2023); Prakash et al.
-(2024) and LMs (i.e., Llama2-7B, Float-7B and Llama3-8B) are applied according to their intended research purpose.
+(2024) and LMs (i.e., Llama2-7B, Float-7B, Llama3-8B, Qwen1.5-7B and Pythia-6.9B) are applied according to their intended research purpose.
 The synthetic datasets we adopted in this work are automatically created by strictly following the rule (or pattern) of the existing dataset, where the entities and attributes are sampled from a pool of wide variety of one-token names and concepts.
 Therefore, there is no ethical concern on human annotation bias and semantic biases.
-The datasets and code will be publicly available to ensure the reproducibility of our experiments.
 Appendix A Appendix
@@
 (2023).
-Different from this line of research, this work studies in-context representations of relations and analyzes how they are stored in model activations.
+Different from this line of research, this work studies in-context representations of relations and analyzes how they are represented in model activations.
 Mechanistic Interpretability Notable progress has been made in uncovering circuits performing various tasks within LMs Elhage et al.
@@
 Feng and Steinhardt (2023) introduce a Binding ID Mechanism for explaining the binding problem, state that LMs use the abstract concept BI to internally mark entity-attribute pairs.
-However, they does not directly capture BI information from activations.
-Therefore, they have not answered how LMs store the BI information into entity activations, how to localize the BI information and whether the localized BI information causally affect the model binding behaviour.
+However, they have not captured the OI information from entity activations that directly determines the binding behaviour.
 A.2 Partial least squares regression and PCA
@@
 (2001), a supervised DR method.
-PLS extracts a set of ordered latent variables that maximizes the co-variability between the features (e.g., activations) and the scores to be predicted (e.g., BI).
+PLS extracts a set of ordered latent variables that maximizes the co-variability between the features (e.g., activations) and the scores to be predicted (e.g., OI).
 We perform PCA and PLS on a development set and compare their regression cures in Figure 11.
-We can observe that both the first PCA component and the first PLS direction contain almost all information about BI of target entity, because their regression score is close to one.
-The consistency indicates that PCA is an effective method to capture BI.
+We can observe that both the first PCA component and the first PLS direction contain almost all information about OI of target entity, because their regression score is close to one.
+The consistency indicates that PCA is an effective method to capture OI subspace.
 Figure 11: Regression curves for PLS and PCA.
-A.3 Layer-wise Intervention
-To determine how well BI subspace from different layers of a LM causally affects the model output, we perform layer-wise BI-PC based intervention mentioned in Section (§4.3) on our development set.
-In Figure 12, we can observe that BI subspace from middle layers (i.e., from layer7 to layer15, especially layer8) significantly affect the computation of binding, and interestingly, these layers also overlap with the ones that clearly encode BI information, as shown in Figure 2.
-Based on such analysis, we select the layer to perform activation patching.
-Figure 12: Average Logit Difference (LD) and logit flip for layer-wise BI-PC based intervention on Llama2-7B, where x axis denotes the layer, the colored zone indicates the layers that are sensitive to the intervention, and the vertical line represents the most sensitive one (i.e., Layer 8), Y axis denotes the average LD and the proportion of inferred attributes (excluding the original one) in Figure 12(a) and Figure 12(b) respectively.
-A.4 Layer-wise Embedding Visualization on Llama3-8B and Float-7B
+A.3 ICA for OI Subspace
+Independent Component Analysis (ICA) is a statistical method to reveal hidden subcomponent from multivariate signal.
+Similar to PCA, ICA is utilized as a dimension reduction method.
+We apply ICA to analyze the activations of entities and visualize the subspace in Figure 12.
+We can observe that the distribution is generally similar to the one of PCA.
+In addition, we also conduct causal intervention (i.e., activation steering) based on the subspace and present the results in Figure 13.
+We can observe that the results are also similar to those from PCA based intervention.
+This indicates that besides PCA, ICA can also be used as a computational method to capture OI subspace.
+Figure 12: Subspace visualization from ICA on Llama2-7B, where “BI” primarily denotes OI.
+Figure 13: Logit Difference and Logit Flip for ICA based activation steering on the entity tracking dataset (i.e., r: is_in) on Llama2-7B.
+A.4 Layer-wise Intervention
+To localize the layer that contributes to binding behaviour, we perform layer-wise OI-PC based intervention mentioned in Section (§4.3) on our development set.
+In Figure 14, we can observe that OI subspace from middle layers (i.e., from layer7 to layer15, especially layer8) significantly affect the computation of binding, and interestingly, these layers also overlap with the ones that clearly encode OI information, as shown in Figure 2.
+Based on the analysis, we select the layer to perform activation patching.
+Figure 14: Average Logit Difference (LD) and logit flip for layer-wise OI-PC based intervention on Llama2-7B, where x axis denotes the layer, the colored zone indicates the layers that are sensitive to the intervention, and the vertical line represents the most effective layer (i.e., Layer 8), Y axis in Figure 14(a) and Figure 14(b) denotes the average LD and the proportion of inferred attributes (excluding the original one) respectively.
+A.5 Layer-wise Embedding Visualization on Llama3-8B and Float-7B
 (a) Llama3-8B
 (b) Float-7B
-Figure 13: Layer-wise BI subspace visualization on Llama3-8B and Float-7B.
-A.5 Layer-wise Embedding Visualization after Masking Binding Information
-The counter argument is that the captured subspace only represents the positional information.
-To test the claim, we attempt to mask the context around the entities and attributes with random two-letter tokens (e.g., “td”) so as to ablate the binding information and keep the positional information.
-For instance, we convert “the document is in Box Q , the bus is in Box F, …” as “pl document td cy wa Q br fl bus ti eq fs F …” so that there is no relational information in the latter.
-Figure 14 compares the BI subspace distribution between with and without binding information.
-We can observe that ablation of binding information tangles the distribution so that there is no clear clustering for each BI (e.g., by comparing layer14).
-This indicates that our discovered subspace encodes binding information.
-(a) w/ Binding Information
-(b) w/o Binding Information
-Figure 14: Layer-wise BI subspace visualization for w/ and w/o binding information on Llama2-7B.
-A.6 Case Study on Llama2-7B
+Figure 15: Layer-wise OI subspace visualization on Llama3-8B and Float-7B.
+A.6 OI Subspace on other LM Families
+(a) Qwen1.5-7B
+(b) Pythia-6.9B
+Figure 16: Layer-wise OI subspace visualization on Qwen1.5-7B and Pythia-6.9B.
+(a) LD for Qwen1.5-7B.
+(b) LF for Qwen1.5-7B.
+(c) LD for Pythia-6.9B.
+(d) LF for Pythia-6.9B.
+Figure 17: Logit Difference (LD) and Logit Flip (LF) for activation patching on the entity tracking dataset (i.e., r: is_in).
+A.7 Case Study on Llama2-7B
 Answer for # Step
@@
 Person Carl is selling the mouse fruit bug grass cookie cookie
-Table 7: Attributes inferred by Llama2-7B as a result of directed activation patching along BI-PC in the BI subspace on the dataset of “r: sell”, where color denotes the BI.
+Table 7: Attributes inferred by Llama2-7B as a result of directed activation patching along OI-PC on the dataset of “r: sell”, where color denotes the BI.
 Answer for # Step
@@
 Person Angel applies the bucket canvas cargo plain floor heavy
-Table 8: Attributes inferred by Llama2-7B as a result of directed activation patching along BI-PC in the BI subspace on the dataset of “r: apply”, where color denotes the BI.
+Table 8: Attributes inferred by Llama2-7B as a result of directed activation patching along OI-PC on the dataset of “r: apply”, where color denotes the BI.
 Answer for # Step
@@
 Person Cole moved the neck pad cone tail toe toe
-Table 9: Attributes inferred by Llama2-7B as a result of directed activation patching along BI-PC in the BI subspace on the dataset of “r: move”, where color denotes the BI.
+Table 9: Attributes inferred by Llama2-7B as a result of directed activation patching along OI-PC on the dataset of “r: move”, where color denotes the BI.
 Answer for # Step
@@
 Person Van brings the note block crown car magnet magnet
-Table 10: Attributes inferred by Llama2-7B as a result of directed activation patching along BI-PC in the BI subspace on the dataset of “r: bring”, where color denotes the BI.
+Table 10: Attributes inferred by Llama2-7B as a result of directed activation patching along OI-PC on the dataset of “r: bring”, where color denotes the BI.
 Answer for # Step
@@
 Person Fran pushes the lever cord vent thumb thumb ear
-Table 11: Attributes inferred by Llama2-7B as a result of directed activation patching along BI-PC in the BI subspace on the dataset of “r: push”, where color denotes the BI.
-A.7 Activation Steering on Llama2-7B
+Table 11: Attributes inferred by Llama2-7B as a result of directed activation patching along OI-PC on the dataset of “r: push”, where color denotes the BI.
+A.8 Activation Steering on Llama2-7B
 (a) r: is_in
@@
 (f) r: push
-Figure 15: Logit Difference (LD) for BI subspace based activation steering across datasets on Llama2-7B, where x axis represents the intervention of $\mathbf{s}_{0\rightarrow bi}$ on the activation of $e_{0}$.
+Figure 18: Logit Difference (LD) for OI subspace based activation steering across datasets on Llama2-7B, where x axis represents the intervention of $\mathbf{s}_{0\rightarrow bi}$ on the activation of $e_{0}$.
 Here, $l=8$ and $\alpha=1.25$.
@@
 (f) r: push
-Figure 16: Logit flip for BI subspace based activation steering across datasets on Llama2-7B, where x axis represents the intervention of $\mathbf{s}_{0\rightarrow bi}$ on the activation of $e_{0}$.
-A.8 Activation Patching on Float-7B
+Figure 19: Logit flip for OI subspace based activation steering across datasets on Llama2-7B, where x axis represents the intervention of $\mathbf{s}_{0\rightarrow bi}$ on the activation of $e_{0}$.
+A.9 Results on Fine-Tuned LM
+Kramár et al.
+(2024); Kim et al.
+(2024) claim that the code fine-tuned LM, such as Float-7B Prakash et al.
+(2024) outperforms the pretrained LM on the entity tracking task Kim and Schuster (2023); Prakash et al.
+(2024).
+Since the code fine-tuned LM performs well on the entity tracking task that requires the OI subspace based computation, we hypothesize that OI subspace also exists in the code fine-tuned LM and the intervention along OI-PC will causally affect the model output.
+To prove the hypothesis, we conduct the intervention on Float-7B and show results in Figure 20 and Figure 21.
+We found that the OI subspace based intervention on Float-7B achieves the similar results as on Llama2-7B, indicating that the OI subspace not only exists in the pretrained LM but also in the fine-tuned one.
+In addition, adding the same step value (i.e., $v$) on Float-7B will achieve higher LD value than Llama2-7B, indicating that the code fine-tuned LM is more sensitive to the OI subspace based intervention.
+For instance, the maximum LD of $a_{4}$ in the former is around $10$, and it is $2$ times larger than the one in the latter, which is around $5$.
+This might partially explains why the code fine-tuned LM performs better than the original one, because code fine-tuning might enhance the function of OI subspace so that it is more sensitive on the intervention and more effective on the in-context entity tracking task.
 (a) r: is_in
@@
 (f) r: push
-Figure 17: Logit Difference (LD) for BI-PC based intervention across datasets on Float-7B, where x axis denotes the number of intervention steps on $e_{0}$, y axis does the LD, BI_i represents each target attribute and the light yellow bottom line indicates the LD of original attribute (i.e., $a_{0}$).
+Figure 20: Logit Difference (LD) for OI-PC based intervention across datasets on Float-7B, where x axis denotes the number of intervention steps on $e_{0}$, y axis does the LD, BI_i represents each target attribute and the light yellow bottom line indicates the LD of original attribute (i.e., $a_{0}$).
 Here, $l=10$, $v=2.55$, and $\alpha=5.0$.
@@
 (f) r: push
-Figure 18: Logit flip for BI-PC based intervention across datasets on Float-7B, where x axis denotes the number of intervention steps on $e_{0}$, y axis does the proportion of each inferred attribute in model output.
-A.9 Activation Patching on Llama3-8B
+Figure 21: Logit flip for OI-PC based intervention across datasets on Float-7B, where x axis denotes the number of intervention steps on $e_{0}$, y axis does the proportion of each inferred attribute in model output.
+A.10 Activation Patching on Llama3-8B
 (a) r: is_in
@@
 (f) r: push
-Figure 19: Logit Difference (LD) for BI-PC based intervention across datasets on Llama3-8B, where x axis denotes the number of intervention steps on $e_{0}$, y axis does the LD, BI_i represents each target attribute and the blue line indicates the LD of original attribute (i.e., $a_{0}$).
+Figure 22: Logit Difference (LD) for OI-PC based intervention across datasets on Llama3-8B, where x axis denotes the number of intervention steps on $e_{0}$, y axis does the LD, BI_i represents each target attribute and the blue line indicates the LD of original attribute (i.e., $a_{0}$).
 Here, $l=10$, $v=0.65$, and $\alpha=2.0$.
@@
 (f) r: push
-Figure 20: Logit flip for BI-PC based intervention across datasets on Llama3-8B, where x axis denotes the number of intervention steps on $e_{0}$, y axis does the proportion of each inferred attribute in model output.
-A.10 Activation Patching on the New Dataset with Interjections
+Figure 23: Logit flip for OI-PC based intervention across datasets on Llama3-8B, where x axis denotes the number of intervention steps on $e_{0}$, y axis does the proportion of each inferred attribute in model output.
+A.11 New Dataset with Interjections
+The dataset is created by inserting a sequence of interjections after the first attribute entity pair, as illustrated in Table 12.
+Since there is no BI information in the interjection (e.g., $j^{p3}$), adding it only changes the PI of its following entities and attributes.
+We set the number of interjections as that the PI of last interjection token is larger than the last PI of its original input (e.g., $pi>p5$ in Table 12).
+Based on this dataset, we conduct the same intervention on its OI subspace, as mentioned in Section (§4.3).
+The counter argument is that the subspace only captures Position Information (PI), and the intervening step only changes the PI information.
+Specifically, adding one unit of $v$ on $\overrightarrow{e_{0}^{p1}}$ might convert the PI of target entity from $p1$ to $p3$, and $p3$ is the PI of $e_{1}$, and its attribute is $a_{1}$, as shown in Table 12, and thus the LM swaps the answer from $a_{0}$ to $a_{1}$.
+If it is true, then the same intervention will not change the answer on the new dataset, because following the counter argument, after adding one unit of $v$ on $\overrightarrow{e_{0}^{p1}}$, the PI of target entity becomes $p3$, and $p3$ is the PI of $j^{p3}$ (i.e., an interjection token).
+The LM thus would not select $a_{1}$ as its answer.
+However, the results on Figure 24 and Figure 25 show that the subspace intervention on the new dataset achieves similar results as the original one, as shown in Figure 3 and Figure 4, proving the counter argument wrong and indicating the independence between OI subspace and PI.
+Input (original)
+$a_{0}^{p0}\,r\,e_{0}^{p1},\,a_{1}^{p2}\,r\,e_{1}^{p3},\,a_{2}^{p4}\,r\,e_{2}^{p5}.\quad e_{0}^{p1}\ r^{-1}\ ?$
+Input (with interjection)
+$a_{0}^{p0}\,r\,e_{0}^{p1},\,j^{p2}\,j\,j^{p3}\,...j^{pi}\,a_{1}^{pi+1}\,r\,e_{1}^{pi+1}...\quad e_{0}^{p1}\ r^{-1}\ ?$
+Table 12: Simplified expression of original input and the one modified with a sequence of interjections, where $j^{p3}$ denotes an interjection $j$, such as “ah”, with position of $p3$, which is also the position of $e_{1}^{p3}$ in the original input.
 (a) r: is_in